### v1

# Dynamic Post-Hoc Neural Ensemblers

Abstract
Ensemble methods are known for enhancing the accuracy and robustness of machine learning models by combining multiple base learners. However, standard approaches like greedy or random ensembles often fall short, as they assume a constant weight across samples for the ensemble members. This can limit expressiveness and hinder performance when aggregating the ensemble predictions. In this study, we explore employing neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively. Motivated by the risk of learning low-diversity ensembles, we propose regularizing the model by randomly dropping base model predictions during the training. We demonstrate this approach lower bounds the diversity within the ensemble, reducing overfitting and improving generalization capabilities. Our experiments showcase that the dynamic neural ensemblers yield competitive results compared to strong baselines in computer vision, natural language processing, and tabular data.
1 Introduction
Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced predictive performance over single-model predictions. Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al., 2017). Combined with their relative simplicity, these properties make ensembling the method of choice for many applications, such as medical imaging and autonomous driving, where reliability is paramount.
Despite these advantages, the process of selecting post-hoc models that are both accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows. Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al., 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation. However, these methods face significant limitations. Specifically, the selection of models to include in the ensemble and the determination of optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data. Although neural networks are good candidates for generating ensembling weights, few studies rely on them as a post-hoc ensembling approach. We believe this happens primarily due to a lack of ensemble-related inductive biases that provide regularization.
In this work, we introduce a novel approach to post-hoc ensembling using neural networks. Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al., 2008). To mitigate the risk of overfitting the validation set, we introduce a regularization technique inspired by the inductive biases inherent to the ensembling task. Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al., 2014).
In summary, our contributions are as follows:
  1. 1.
    We propose a simple yet effective post-hoc ensembling method based on a neural network that dynamically ensembles base models.
  2. 2.
    To prevent the formation of low-diversity ensembles, we introduce a regularization technique that involves randomly dropping base model predictions during training. We demonstrate theoretically that this lower bounds the diversity of the generated ensemble. Additionally, we ablate its effect through various experiments.
  3. 3.
    Through extensive experiments, we show that Neural Ensemblers consistently select competitive ensembles across a wide range of data modalities, including tabular data (for both classification and regression), computer vision, and natural language processing.
To promote reproducibility, we have made our code publicly available in the following anonymous repository  [^1] https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We hope that our codebase, along with the diverse set of benchmarks used in our experiments, will serve as a valuable resource for the development and evaluation of future post-hoc ensembling methods.
2 Background and Motivation
Post-Hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$. These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatination of the base models predictions. While the base models are estimated using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
$\theta\in\operatorname*{arg\,min}_{\theta}\;\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}\mathcal{L}(f(z(x);\theta),y).$ (1)
In the general case, this objective function can be optimized using gradient-free optimization method such as evolutionary algorithms (Purucker & Beel, 2023b) or greedy search (Caruana et al., 2004) together with a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
$f(z(x);\theta)=\sum_{m}\theta_{m}z_{m}(x).$ (2)
Additionally, if we constraint the ensembler weights such that $\forall_{i}\theta_{i}\in\mathbb{R}_{+}$ and $\sum_{i}\theta_{i}=1$ and assume probabilistic base models $z_{m}(x)=p(y|x,m)$, then we can interpret Equation 2 as:
$p(y|x)=\sum_{i}p(y|x,m)p(m),$ (3)
which is referred to as Bayesian Model Average, and uses $\theta_{m}=p(m)$. In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacker model parametrized by $\beta$.
2.1 Motivating Dynamic Ensembling
Figure 1: Wrong Models Per Samples Across Meta-Datasets. Every dark cell represents data instances where a model’s prediction is wrong. Different models fail on different instances, therefore, only instance-specific dynamic ensembles are optimal.
We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental metadatasets. In reality, Equation 3 does not specify the distribution $p(m)$. Generally, it is safe to assume that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we dynamically select the optimal aggregation $\theta_{m}(x)=p(m|m)$ on a per-data point basis, instead of a static $\theta$. To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas & Erickson, 2023)), QuickTune (Computer Vision (Arango et al., 2024)), FTC (NLP-Section D.1) and NasBench 201 (NAS for Computer Vision (Dong & Yang, 2020)). Then, we compute the per-sample error for 100 models in 100 samples. We report the results in Figure 1, indicating failed predictions with dark colors. We observe that models make different errors across samples, demonstrating the lack of optimality for static ensembling weights.
3 Neural Ensemblers
Figure 2: Architecture of Neural Ensemblers (classification). The stacking mode uses a single MLP shared across base model class predictions. It outputs the logit per class, used for computing the final probability via SoftMax. In Model Averaging mode, it generates the unnormalized weights for every model, which are normalized with SoftMax.
We use neural networks as ensemblers by training on the base model predictions $z(x)=[z_{1}(x);...;z_{M}(x)]$ as input. For regression, $z(x):\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs thebase models’ point predictions given by $x\in\mathbb{R}^{D}$, while for classification $z(x;c):\mathbb{R}^{D}\rightarrow[0,1]^{M}$ returns the probabilities predicted by the base models for class $c$. In our discussion we consider two functional modes for the ensemblers: as network outputting weights for model averaging or as a direct stacking model that outputs the prediction. In stacking mode for regression, we aggregate the base model point predictions using a neural network to estimate the final prediction $\hat{y}=\phi(z;\beta)$, where $\beta$ are the network parameters. In the model-averaging mode, the Neural Ensembler outputs the weights $\theta_{m}(z;\beta)$ to combine the model predictions as in Equation 4.
$\hat{y}=\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}(x).$ (4)
Regardless of the functional mode, the Neural Ensembler has a different output $\hat{y}$ for regression and classification. In regression, the output $\hat{y}$ is a point estimation of the mean for a normal distribution such that $p(y|x;\beta)=\mathcal{N}(\hat{y},\sigma)$. For classification, the input is the probabilistic prediction of the base models per class $z_{m}(x;c)=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z(x),\beta)$. We optimize $\beta$ by minimizing the negative log-likelihood over the validation dataset $\mathcal{D}_{\mathrm{Val}}$ as:
$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta).$ (5)
3.1 Architecture
We discuss the architectural implementation of the Neural Ensembler for the classification case, which we show in Figure 2. For the stacking mode, we use an MLP that outputs the logit $\phi(z(x;c);\beta):\mathbb{R}^{M}\rightarrow[0,1]$ of each class $c$. Specifically, the network receives as inputs the base models’ predictions $z(x,c)$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e. $\phi(z;c)$. The model predictions per class are fed independently, enabling sharing the network parameters $\beta$. Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z;c)}}{\sum_{c^{\prime}}\exp^{\phi(z;c^{\prime})}}$, with $\mathrm{MLP}(z(x;c);\beta)$. In regression, the final prediction is the output $\phi(z(x);\beta)$.
For model averaging mode, we use a novel architecture based on a Deep Set (Zaheer et al., 2017) embedding of the base models predictions. We compute the dynamic weights $\theta_{m}(z;\beta)=\frac{\exp f_{m}(z;\beta)}{\sum_{m^{\prime}}f_{m^{\prime}}(z;\beta)}$, where the unnormalized weight per model $f_{m}(z;\beta)$ is determined via two MLPs. The first one $\text{MLP}_{1}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ embeds the predictions per class $z(x,c^{\prime})$ into a latent dimension of size $H$, whereas the second network $\text{MLP}_{2}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ aggregates the embeddings and outputs the unnormalized weights, as shown in Equation 6. Notice that the Neural Ensemblers’ input dimension and number of parameters are independent of the number of classes, due to our proposed parameter-sharing mechanism.
$f(z;m,\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z(x;c^{\prime});\beta_{1}\right);\beta_{2}\right).$ (6)
3.2 The Risk of Diversity Collapse
In this section, we elaborate on one risk that might arise during learning ensembles, as a base model might be highly correlated with the target in the validation set, when building ensembles. We dub this model a Preferred Base Model.
Definition 1 (Preferred Base Model).
Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$. $z_{p}$ is the Preferred Base Model if it has the highest sample correlation to the target, i.e. $\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
Proposition 1.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{p}\}$, given the conditions of Definition 1.
Sketch of Proof.
Given that the base models predictions are uncorrelated, their correlations $\rho_{z_{m},z_{p}}=0,\forall m\neq p$, then it holds that $-\sqrt{1-\rho_{z_{p},y}^{2}}\leq\rho_{z_{m},y}\leq\sqrt{1-\rho_{z_{p},y}^{2}},\forall z_{m}\in\mathcal{F}/\{z_{p}\}$. We provide more details in the Appendix A. ∎
On the other hand, an important aspect when building ensembles is guaranteeing diversity among the models (Wood et al., 2023; Jeffares et al., 2024). This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al., 2022; Purucker et al., 2023). A common way to measure diversity is the ambiguity (Krogh & Vedelsby, 1994), which can be derived after decomposing the loss function (Jeffares et al., 2024). Unfortunately, even simple ensembles risk obtaining low diversity although their base models are uncorrelated. As we will shortly show, this happens especially when there is a preferred base model.
Proposition 2 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$, i.e. $\lim_{\rho_{z_{p},y}\to 1}\alpha=0$.
We give a detailed proof in the Appendix A. However, we can build the intuition after noticing that given Proposition 1, $\lim_{\rho_{z_{p},y}\to 1}\bar{z}=z_{p}$.
3.3 Base Models’ DropOut
Input: Base model predictions $\{z_{1}(x),...,z_{M}(x)\}$, validation data $\mathcal{D}_{\mathrm{Val}}$, probability of retaining $\gamma$, mode $\in\{\mathrm{Stacking,Averaging}\}$.
Output: Neural Ensembler’s parameters $\beta$
1
2Initialize randomly parameters $\beta$ ;
3
4while $done$ do
5
6      Sample masking vector $r\in\mathbb{R}^{M},r_{i}\sim\mathrm{Ber}(\gamma)$;
7
8      Mask base models predictions $z_{\mathrm{drop}}(x)=r\odot z(x)$ ;
9
[…]
11             Compute predictions $\hat{y}=\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}};\beta\right)$ ;
12      else
13
14            Compute weights $\theta\left(\frac{1}{\gamma}z_{\mathrm{drop}};m,\beta,r\right)$ using Equation 7 ;
15
16            Compute predictions $\hat{y}$ using Equation 4 ;
17
18       end if
19      Update neural ensembles parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
20 end while
21
22return $\beta$;
23
Algorithm 1 Training Algorithm for Neural Ensemblers with Base Models’ DropOut
Using Neural Ensemblers tackles the need for dynamical ensembling. Moreover, it gives additional expressivity associated with neural networks. However, there is also a risk of overfitting and diversity collapse. As we showed in Section 3.2, this happens when there is a preferred model on which the ensembler mainly relies while neglecting other base model predictions. Although it might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples. Inspired by previous work (Srivastava et al., 2014), we propose to drop some base models during training forward passes. Intuitively, this forces the ensembler to rely on different base models to perform the predictions, instead of merely using the preferred base model(s).
Formally, we mask the inputs such as $r_{m}\cdot z_{m}(x)$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate. We also mask the weights when using model averaging:
$\theta(z;m,\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
As DropOut changes the scale of the inputs, the weight scaling rule should be applied during inference by multiplying the dropped variables by the retention probability $\gamma$. Alternatively, we can scale the variables during training by $\frac{1}{\gamma}$. In Algorithm 1, we detail how to train the Neural Ensemblers by dropping base model predictions. It has two modes, acting as a direct stacker or as a model averaging ensembler. We demonstrate that base models’ DropOut avoids diversity collapse by upper bounding the diversity as stated in Proposition 6, even for the simplest ensembling case.
Proposition 3 (Avoiding Diversity Collapse).
As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
Sketch of Proof. We want to compute $\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$. By using $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$, and assuming, without loss of generality, that the predictions are standardized, we obtain $\mathbb{V}(r\cdot z_{m})=\gamma$. This lead as to $\lim_{\rho_{z_{p},y}\to 1}\alpha=1-\gamma$, after following a procedure similar to Proposition 2. We provide the complete proof in Appendix A.
4 Experiments and Results
4.1 Proof-of-Concept: Ensemble of Quadratic Functions
To further highlight the importance of the dynamical ensemble with diverse base models, we propose a simple regression problem with a third-degree polynomial with the ground truth function $y_{\mathrm{true}}=1.33x^{3}-0.77x^{2}-0.31x-1$, where $x\in\left[-1,1\right]$. We fit three second-degree polynomials on different subsets of the training dataset, to obtain three base functions $z_{1}(x),z_{2}(x),z_{3}(x)$. As ensemblers we consider i) a static model average $\sum_{i}\theta_{m}z_{m}(x)$, ii) a dynamic model stacker $\phi(z(x);\beta)$ and iii) model average $\sum_{m}\theta_{m}(z(x);\beta)\cdot z_{m}(x)$. In this experiment, $\phi(\cdot)$ and $\theta_{m}(\cdot)$ are two-layer MLPs with 10 hidden neurons with parameters $\beta$ and trained on the validation data. Figure 3 shows the specific data, base functions, and learned ensemblers. The rightmost plot shows that the dynamic model averaging (MSE=$0.0101$) and stacker (MSE=$0.0055$), by expressing more complex functions, better model the ground truth than the static ensemble (MSE=$0.1662$).
Figure 3: Intuition Example for Dynamical Ensembling. When considering three base models $z_{1}(x),z_{2}(x),z_{3}(x)$, the dynamical ensemblers $\phi(\cdot)$ and $\theta_{m}(\cdot)$ depending on $x$ achieve better performance than a static model average, as observable in the right-most figure.
4.2 Experimental Setup
Meta-Datasets.
In our experiments, we utilize four meta-datasets with pre-computed predictions, which allows us to simulate ensembles without the need to fit models. These meta-datasets cover diverse data modalities, including Computer Vision, Tabular Data, and Natural Language Processing. Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that require model fitting. Table 1 reports the main information related to these datasets. Particularly for Nasbench, we created 2 versions by subsampling 100 and 1000 models. The metadataset for Finetuning Text Classifiers (FTC) was generated by ourselves to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al., 2019), Bert (Devlin et al., 2018) and Bart (Lewis et al., 2019). We also generate a set of fitted Scikit-Learn Pipelines on classification datasets. In this case, we stored the pipeline in memory, allowing us to evaluate our method in practical scenarios where the user has fitted models instead of predictions. We detailed information about the creation of these two metadatasets in Appendix D. Information about each dataset lies in the respective referred work under the column Task Information in Table 1.
Table 1: Metadatasets Information.
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
Baselines.
We compare the Neural Ensemblers (NE) with other common and competitive ensemble approaches. 1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al., 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker & Beel, 2023b) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3. We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al., 2022), and 9) Ensemble Optimization (EO) (Levesque et al., 2016). Finally, we report results by using common ML models as stackers, such as 10) SVM, 11) Random Forest, 12) Gradient Boosting, and 13) Logistic/Linear Regression. We used the default configurations provided by Sckit-learn (Pedregosa et al., 2011) for these stackers. The input to the models is the concatenation of all the base models’ predictions. We concatenated the probabilistic predictions from all the classes in the classification tasks. This sometimes produced a large dimensional input space. Finally, we include models from Dynamic Ensemble Search (DES) literature that are in the DESlib library (Cruz et al., 2020) such as KNOP (Cavalin et al., 2013), KNORAE (Ko et al., 2008) and MetaDES (Cruz et al., 2015).
Neural Ensemblers’ Setup.
We train the neural networks for $10000$ update steps, with a batch size of $2048$. If the GPU memory is not enough for fitting the network because of the number of base models, or the number of classes, we decreased the batch size to $256$. Additionally, we used the Adam optimizer and a network with four layers, $32$ neurons, and a probability of keeping base models $\gamma=0.25$, or alternatively a DropOut rate $\delta=0.75$. Notice that the architecture of the ensemblers slightly varies depending on the mode (Stacking or Model Average). For the Stacking mode, we use an MLP with four layers and $32$ neurons with ReLU activations. For MA mode, we use two MLPs as in Equation 6: 1) $\textrm{MLP}_{1}$ has $3$ layers with $32$ neurons, while 2) $\textrm{MLP}_{2}$ has one layer with the same number of neurons. Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune metadataset (extended version).
4.3 Research Questions and Associated Experiments
Table 2: Average Normalized Error.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Class (AUC)
Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random 1.3377±0.2771 0.7283±0.0752 0.7491±0.2480 6.6791±3.4638 4.7284±2.9463 1.4917±1.6980 1.7301±1.8127
Top5 0.9511±0.0364 0.6979±0.0375 0.6296±0.1382 0.6828±0.3450 0.8030±0.2909 0.9998±0.1233 0.9271±0.2160
Top50 1.1012±0.1722 0.6347±0.0395 0.5650±0.1587 1.0662±0.9342 1.0721±0.4671 0.9800±0.1773 0.9297±0.2272
Quick 0.9494±0.0371 0.6524±0.0436 0.5787±0.1510 0.7575±0.2924 0.7879±0.2623 0.9869±0.1667 0.9054±0.2232
Greedy 0.9494±0.0374 0.7400±0.1131 1.0000±0.0000 0.9863±0.4286 0.9297±0.1435 0.9891±0.1693 0.9090±0.2197
CMAES 0.9489±0.0392 0.6401±0.0343 0.5797±0.1575 1.0319±0.5000 0.9086±0.1121 0.9935±0.1953 1.1878±1.1457
Random Forest 0.9513±0.0359 0.6649±0.0427 0.6891±0.3039 1.4738±1.3510 1.2530±0.4875 1.0041±0.2330 1.0924±0.6284
Gradient Boosting 1.0097±0.1033 1.2941±0.5094 1.2037±0.3528 0.8514±0.5003 1.6121±1.7023 1.0452±0.3808 1.0663±0.4884
SVM 0.9453±0.0383 0.6571±0.0483 0.7015±0.3067 1.1921±0.8266 1.4579±0.6233 0.9585±0.2160 1.4701±1.3486
Linear 0.9609±0.0347 0.7891±0.1978 0.7782±0.1941 0.7333±0.4457 0.9291±0.3580 0.9776±0.2844 1.0329±0.4022
MA 1.0709±0.0845 0.6381±0.0349 0.5610±0.1490 1.1548±0.8465 1.2173±0.6107 1.0917±1.0135 0.9977±0.2278
DivBO 1.0155±0.1452 0.6915±0.0536 0.9120±0.1524 1.3935±1.4316 1.0635±0.7587 1.0908±1.0104 1.0899±1.0297
EO 1.0208±0.1159 0.6365±0.0445 0.5704±0.1619 1.0185±0.6464 1.0367±0.4394 1.0851±1.0136 0.9377±0.2310
NE-Stack (Ours) 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
NE-MA (Ours) 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
RQ 1: Can Neural Ensemblers outperform other common and competitive ensembling methods across data modalities?
Figure 4: Results on Scikit-Learn Pipelines.
Experimental Protocol. To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the metadatasets. We run every ensembling method three times for every dataset. In all the methods we use the validation data for fitting the ensemble, while we report the results on the test split. Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and error. For the tabular classification, we compute the ROC-AUC. As these metrics vary for every dataset, we normalize metrics by dividing them by the single-best metric. Therefore, a method with a normalized metric below one is improving on top of using the single best base model. We report the standard deviation across the experiments per dataset and highlight in bold the best method.
Results. The results reported in Table 2 and Table 3 show that our proposed neural networks are competitive post-hoc ensemblers. In general, we observe that the Neural Ensemblers variants obtain either the best (in bold) or second best (italic) performance across almost all meta-datasets and metrics. Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class. This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al., 2020). We hypothesize that dynamic ensembling contributes partially to the strong results for the Neural Ensemblers. However, the expressivity gained is not enough, because it can lead to overfitting. To understand this, we compare to Dynamic Ensemble Selection (DES) methods. Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines metadataset, as we can easily access the fitted models. We report the results of the test split in Figure 4, where we distinguish among four types of models to facilitate the reading: Neural, DES, Stacking and Constant. We can see that Neural Ensemblers are the most competitive approaches, especially on stacking mode. Additionally, we report the metrics on the validation split in Figure 8 (Appendix E), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust in this sense.
Table 3: Average Normalized NLL.
FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random 1.5450±0.5289 0.6591±0.2480 0.7570±0.2900 6.8911±3.1781 5.8577±3.2546 1.7225±1.9645 1.8319±2.1395
Top5 0.8406±0.0723 0.6659±0.1726 0.6789±0.3049 1.5449±1.8358 1.1496±0.3684 1.0307±0.5732 0.9939±0.0517
Top50 0.8250±0.1139 0.5849±0.2039 0.6487±0.3152 3.3068±2.6197 3.0618±2.2960 1.0929±1.0198 1.0327±0.2032
Quick 0.7273±0.0765 0.5957±0.1940 0.6497±0.3030 1.1976±1.1032 0.9747±0.2082 0.9860±0.2201 1.0211±0.1405
Greedy 0.6943±0.0732 0.5785±0.1972 1.0000±0.0000 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
CMAES 1.2356±0.5295 1.0000±0.0000 1.0000±0.0000 4.1728±2.8724 4.6474±3.0180 1.3487±1.3390 1.0281±0.1977
Random Forest 0.7496±0.0940 0.8961±0.3159 0.9340±0.4262 3.7033±2.8145 2.2938±2.2068 1.2655±0.4692 1.0030±0.0871
Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2764±0.4787 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
SVM 0.7990±0.0909 0.7744±0.2967 0.9358±0.5706 5.4377±3.3807 4.0019±3.6601 1.3884±1.4276 2.7975±3.0219
Linear 0.7555±0.0898 0.7400±0.2827 0.8071±0.2206 1.3960±1.2334 1.1031±0.7038 1.1976±1.1024 3.1488±3.2813
MA 0.9067±0.1809 0.5970±0.2034 0.6530±0.3028 4.7921±3.0780 4.0168±2.8560 1.4724±1.9401 1.3342±1.3515
DivBO 0.7695±0.1195 0.7307±0.3061 0.7125±0.3982 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
EO 0.7535±0.1156 0.5801±0.2051 0.6911±0.2875 1.3702±1.6389 0.9649±0.2980 1.0979±1.0289 1.0183±0.0993
NE-Stack (Ours) 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
NE-MA (Ours) 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
RQ 2: Do Neural Ensemblers need a strong group of base models, i.e. found using Bayesian Optimization?
Experimental Protocol. Practitioners use some methods such as greedy ensembling as post-hoc ensemblers, i.e., they consider a set of models selected by a search algorithm such as Bayesian Optimization as base learners. DivBO enhances the Bayesian Optimization by accounting for the diversity in the ensemble in the acquisition function. We run experiments to understand whether the Neural Ensemblers’ performance depends on a strong subset of $50$ base models selected by DivBO, and whether it can help other methods. We conduct additional experiments by randomly selecting $50$ models to understand the impact and significance of merely using a smaller set of base models. We normalize the base of the metric on the single-best base model from the complete set contained in the respective dataset.
Results. We report in Table 4 the results with the two selection methods (random and DivBO) using a subset of common baselines, where we normalize using the metric of the single-best from the whole set of models. We limit the number of baselines for brevity, and the extended version includes more baselines in Table 15. We also include results in separate tables for the two selection methods (Appendix E). We can compare directly with the results in Table 3. We observe that reducing the number of base models with DivBO negatively affects the performance of the Neural Ensemblers. Surprisingly, randomly selecting the subset of base models improves the results in two metadatasets (TR-Class and NB-1000). We hypothesize that decreasing the number of base models is beneficial for these metadatasets. With over $1000$ base models available, the likelihood of identifying a preferred model and overfitting the validation data increases in these metadatasets. Naturally, decreasing the number of base models can also be detrimental for the Neural Ensemblers, as this happens for some metadatasets such as TR-Reg and QT-Micro. In contrast to the Neural Ensemblers, selecting a subset of strong models with DivBO improves the performance for some baselines such as Model Averaging (MA) or TopK ($K=25$). In other words, it works as a preprocessing method for these ensembling approaches. Overall, the results in Tables 4 and 15 demonstrate that Neural Ensemblers do not need a strong group of base models to achieve competitive results.
Table 4: Average Normalized NLL with a Subset of Base Models.
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
NE-Stack (Ours) Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA (Ours) Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
RQ 3: What is the impact of the DropOut regularization scheme?
Experimental Protocol. Firstly, we do a small experiment to observe how the DropOut rate affects the weights when using Neural Ensemblers in MA mode. For this, we use QT-Micro datasets, and save each model’s weights $\theta_{m}$ and error after training the neural network. Subsequently, to understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$. We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset. Therefore, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
Results. We show how the mean of the weights per model is related to the mean error of the models for different DropOut rates in Figure 5. When there is no DropOut some weights are close to one, i.e. they are preferred models. As we increase the value, many models with high weights decrease. If the rate is very high (e.g. $0.9$), we will have many models contributing to the ensemble, with weights different from zero. Our ablation study demonstrates that non-existing or high DropOut are detrimental to the Neural Ensembler performance in general. As shown in Figure 6, this behavior is consistent in all datasets and both modes, but TR-Reg metadataset on Stacking mode. In general, we observe that Neural Ensemblers obtain better performance when using base models’ DropOut.
Figure 5: Mean weights assigned to the base models decrease with DropOut rate. Every data point is a model. The errors and weights are the mean values across many datasets for every model.
Figure 6: Ablation of the DropOut rate.
5 Related Work
Ensembles for Tabular Data. For tabular data, ensembles are known to perform better than individual models (Sagi & Rokach, 2018; Salinas & Erickson, 2023).Therefore, ensembles are often used in real-world applications (Dong et al., 2020), to win competitions (Koren, 2009; Kaggle, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker & Beel, 2023b; Purucker et al., 2023). Methods like Bagging (Breiman, 1996) or Boosting (Freund et al., 1996) are often used to boost the performance of individual models. In contrast, post-hoc ensembling (Shen et al., 2022; Purucker & Beel, 2023a) aggregates the predictions of an arbitrary set of fitted base models. Post-hoc ensembles are build by stacking (Wolpert, 1992; Van der Laan et al., 2007), ensemble selection (a.k.a. pruning) (Caruana et al., 2004; Tsoumakas et al., 2009), dynamic ensemble selection (Ko et al., 2008; Britto Jr et al., 2014), or through a systematic search for an optimal ensemble (Levesque et al., 2016; Shen et al., 2022; Poduval et al., 2024).
Ensembles for Deep Learning. Ensembles of neural networks (Hansen & Salamon, 1990; Krogh & Vedelsby, 1994; Dietterich, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation. Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al., 2017) being the most popular one, which involve independently training multiple initializations of the same network. Extensive empirical studies (Ovadia et al., 2019; Gustafsson et al., 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al., 2015; Gal & Ghahramani, 2016; Welling & Teh, 2011).
Mixture-of-Experts. Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017), where one network is trained with specialized sub-modules that are activated based on the input data. Batch ensembles (Wen et al., 2020; Wenzel et al., 2020) are also closely related, as they aim to capture diverse model behaviors by simultaneously training multiple models with shared weights but different random projections. In contrast to MoE and batch ensembles, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches. Ours additionally considers a regularization strategy that leads to provably more diverse ensembles.
Dynamic Ensemble Selection. Our Neural Ensembler is highly related to dynamic ensemble selection. Both dynamically aggregate the predictions of base models per instance (Cavalin et al., 2013; Ko et al., 2008). Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al., 2018; 2020). In contrast, we use an end-to-end trained neural network to select and weight the base models per instance.
6 Conclusions
In this work, we tackled the challenge of post-hoc ensemble selection and the associated risk of overfitting on the validation set. We introduced the Neural Ensembler, a neural network that dynamically assigns weights to base models on a per-instance basis. To reduce overfitting, we proposed a regularization technique that randomly drops base models during training, which we theoretically showed enhances ensemble diversity. Our empirical results demonstrated that Neural Ensemblers consistently form competitive ensembles across diverse data modalities, including tabular data (classification and regression), computer vision, and natural language processing. In the future, we aim to explore in-context learning (Brown et al., 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input. We discuss broader impact and limitations in Appendix B.
Reproducibility Statement
All code and datasets used in this paper are publicly available at https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We provide the code for running all the experiments, including baselines on all meta-datasets. We also provide instructions for the environment setup to ensure reproducibility. We will deanonymize the link upon acceptance. We further provide all proofs in Appendix A. Moreover, we ensured reproducibility by testing our approach on many meta-datasets and across several domains, for which details can be found in Appendix D. We also compared our method to many baselines commonly used in ensembling, described in Section 4.3. Lastly, we detail additional results and extended data tables in Appendix E.
Acknowledgments
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project v101be. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828 and 499552394 - SFB 1597; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix
Appendix A Proofs
Definition 2 (Preferred Base Model).
Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$. $z_{p}$ is the Preferred Base Model if it has the highest sample correlation with respect the target, i.e. $\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
Proposition 4.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{k}\}$, given the conditions of Definition 1.
Proof.
Let $\rho_{z_{p},z_{m}},\rho_{z_{m},y},\rho_{z_{p},y}$ be the correlations among the preferred model $z_{p}$, a non-preferred model $z_{m}\neq z_{m}$ and the target $y$, then the correlation matrix is given by,
$C=\begin{pmatrix}1&\rho_{z_{p},z_{m}}&\rho_{z_{p},y}\\ \rho_{z_{p},z_{m}}&1&\rho_{z_{m},y}\\ \rho_{z_{p},y}&\rho_{z_{m},y}&1\\ \end{pmatrix}$ (8)
The determinant of the correlation matrix is given by,
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},z_{m}}^{2}-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}+2\rho_{z_{p},z_{m}}\rho_{z_{p},y}\rho_{z_{m},y}$ (9)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}$ (10)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}\geq 0$ (11)
$\rho_{z_{m},y}$ $\leq\sqrt{1-\rho_{z_{p},y}^{2}},$ (12)
where we used the fact that $\rho_{z_{p},z_{m}}=0$, in Equation 10 and that the correlation matrix should be semidefinite in Equation 11, thus its determinant should hold $\mathrm{Det}(C)\geq 0$. This means that $\epsilon=\sqrt{1-\rho_{z_{p},y}^{2}}$ is an upper bound for the absolute value of the correlation $\rho_{z_{m},y}$. The limit of the upper bound $\epsilon$ as $\rho_{z_{p},y}\to 1$ is 0 and thus $\rho_{z_{m},y}=0$,
$\lim_{\rho_{z_{p},y}\to 1}$ $\epsilon=\lim_{\rho_{z_{p},y}\to 1}\sqrt{1-\rho_{z_{p},y}^{2}}=0$ (13)
$\lim_{\rho_{z_{p},y}\to 1}$ $\rho_{z_{m},y}=0.$ (14)
Since $\rho_{z_{m},y}$ is bounded by $-\epsilon\leq\rho_{z_{m},y}\leq\epsilon$.
Proposition 5 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$.
Proof.
Firstly, we show that under conditions of Definition 1, $\theta_{m}=\hat{\rho}_{z_{m},y}$, where $\hat{\rho}_{z_{m},y}$ is the sample correlation. Assume that we have a matrix of evaluations $X\in\mathrm{R}^{N\times M}$, with components $x_{n,m}$ indicating the prediction of the $m$-th model for the $n$-th instance. If $Y\in\mathbb{R}^{N\times 1}$ is the ground-truth, then the ensemble weights $\Theta\in\mathbb{R}^{M\times 1}$ can be computed with closed form solution of the objective $\min_{\Theta}||Y-X\Theta||^{2}$ as: $\Theta=(X^{T}X)^{-1}X^{T}Y$.
Without loss of generality, we assume that the random variables are standardized. Given that every row of $X$ have samples from uncorrelated random variables, then $X^{T}X\approx I$, denoting $I$ as the identity matrix, thus
$\Theta=X^{T}Y$ (15)
Every component of $\Theta$ can expressed as summatorias $\theta_{m}=\sum_{n}x_{n,m}y_{n}$. Note that $x_{n,m}$ is a standardized sample (i.e. mean 0, variance 1) from the random variable $z_{m}$, thus it equals the sample correlation $\theta_{m}=\sum_{n}x_{n,m}y_{n}=\hat{\rho}_{z_{m},y}$.
Now we develop the formal demonstration of Proposition 2. We note to the right side a hint to the criteria that was applied to derive every equation.
$\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}$ $\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$ Definition of $\alpha$. (16)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{E}\left[(z_{m}-\bar{z})^{2}\right]$ Properties of expectations. (17)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{E}[z_{m}-\bar{z}]^{2}+\mathbb{V}[z_{m}-\bar{z}]\right)$ Expectation of a squared variable. (18)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{V}[z_{m}-\bar{z}]$ Means of $z_{m}$ and $\bar{z}$ are 0. (19)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{V}[z_{m}]-\mathbb{V}[\bar{z}]\right)$ Property of sum of variances. (20)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}[\bar{z}]\right)$ Variance of $z_{m}$ is 1. (21)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}\left[\sum_{m^{\prime}}\theta_{m^{\prime}}z_{m^{\prime}}\right]\right)$ Definition of $\bar{z}$. (22)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\mathbb{V}\left[z_{m^{\prime}}\right]\right)$ Properties of variance. (23)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Variance of $z_{m}$ is 1. (24)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Property of sum of limits. (25)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\cdot\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Product-sum property of limits. (26)
$=\left[\sum_{m}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m},y}\right]$ $\cdot\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given that $\theta_{m}=\hat{\rho}_{z_{m},y}\approx\rho_{z_{m},y}$. (27)
=1=1$\cdot$= 1 ⋅ $\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given Proposition 1. (28)
$=1-$ $\sum_{m^{\prime}}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m^{\prime}},y}^{2}$ Properties of limits. (29)
$\lim_{\rho_{z_{p},y}\to 1}$ $\alpha=0$ Given Proposition 1. (30)
Lemma 1.
When using dropout, it still holds that $\theta_{m}=\hat{\rho}_{z_{m},y}.$
Proof.
$\min_{\Theta}\mathbb{E}_{R\sim\mathrm{Ber(\gamma)}}||Y-(R*X)\Theta||^{2}$ (31)
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Gamma\Theta||^{2},$ (32)
where $\Gamma=(\mathrm{diag(X^{T}X)})^{1/2}$. If we use the uncorrelation of the base models, we obtain:
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2}$ (33)
$=\min_{\Theta}||Y-\tilde{X}\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2},$ (34)
which is equivalent to ridge regression, with closed form solution:
$\Theta=$ $\left({\tilde{X}}^{T}{\tilde{X}}+\gamma(1-\gamma)I\right)^{-1}{\tilde{X}}^{T}Y$ (35)
$=$ $\left([\gamma^{2}+\gamma(1-\gamma)]I\right)^{-1}{\tilde{X}}^{T}Y$ (36)
$=$ $\left(\gamma I\right)^{-1}{\tilde{X}}^{T}Y$ (37)
$=$ $\frac{1}{\gamma}{\tilde{X}}^{T}Y$ (38)
$=$ $X^{T}Y.$ (39)
Proposition 6 (Avoiding Diversity Collapse).
As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
Proof.
We follow a similar procedure as for Proposition 2, by considering $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$. We demonstrate that $\mathbb{V}(r\cdot z_{m})=\gamma$, given that $r\sim\mathrm{Bernoulli}(\gamma)$.
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})\cdot\mathbb{V}(z_{m})+\mathbb{V}(z_{m})\cdot\mathbb{E}(r_{m})^{2}+\mathbb{V}(r_{m})\cdot\mathbb{E}(z_{m})^{2}$ (40)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})+\mathbb{E}(r_{m})^{2}$ (41)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma(1-\gamma)+\gamma^{2}$ (42)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma.$ (43)
Then, we evaluate the variance of the ensemble using DropOut $\mathbb{V}(\bar{z})$:
$\mathbb{V}(\bar{z})=\mathbb{V}{\left(\sum_{m}r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (44)
$\mathbb{V}(\bar{z})=\sum_{m}\mathbb{V}{\left(r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (45)
$\mathbb{V}(\bar{z})=\sum_{m}\theta^{2}_{m}\mathbb{V}{\left(r_{m}\cdot z_{m}\right)}$ (46)
$\mathbb{V}(\bar{z})=\gamma\sum_{m}\theta^{2}_{m}.$ (47)
Applying Equation 47 into Equation 22, we obtain:
$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\gamma_{m}\cdot\theta_{m}(z_{m}-\bar{z})^{2}\right]$ (48)
$=\lim_{\rho_{z_{p},y}\to 1}\sum_{m}\theta_{m}\left(1-\gamma\sum_{m}\theta^{2}_{m}\right)$ (49)
$=\lim_{\rho_{z_{p},y}\to 1}\left(1-\gamma\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)$ (50)
$=1-\gamma\cdot\lim_{\rho_{z_{p},y}\to 1}\left(\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)$ (51)
$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=1-\gamma.$ (52)
Appendix B Limitations and Broader Impact.
While our proposed method offers several advantages for post-hoc ensemble selection, it is important to recognize its limitations. Unlike simpler ensembling heuristics, our approach requires tuning multiple training and architectural hyperparameters. Although we employed a fixed set of hyperparameters across all modalities and tasks in our experiments, this robustness may not generalize to all new tasks. In such cases, hyperparameter optimization may be necessary to achieve optimal performance. However, this could also enhance the results presented in this paper. Our approach is highly versatile and can be seamlessly integrated into a wide variety of ensemble-based learning systems, significantly enhancing their predictive capabilities. Because our method is agnostic to both the task and modality, we do not expect any inherent negative societal impacts. Instead, its effects will largely depend on how it is applied within different contexts and domains, making its societal implications contingent on the specific use case.
Appendix C Related Work Addendum
Ensemble Search via Bayesian Optimization. Ensembles of models with different hyperparameters can be built using Bayesian optimization by iteratively swapping a model inside an ensemble with another one that maximizes the expected improvement (Levesque et al., 2016). DivBO (Shen et al., 2022) and subsequent work (Poduval et al., 2024) combine the ensemble’s performance and diversity as a measure for expected improvement. Besides Bayesian Optimization, an evolutionary search can find robust ensembles of deep learning models (Zaidi et al., 2021). Although these approaches find optimal ensembles, they can overfit the validation data used for fitting if run for many iterations.
Appendix D Details On Metadatasets
We provide general information about the datasets in Table 5.
Table 5: Metadatasets Information
Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
D.1 Finetuning Text Classifiers (FTC) Metadataset
Table 6: Search Space.
Hyperparameter Values
Model GPT2, Bert-Large, Albert-Large, Bart-Large, T5-Large
Learning Rate 0.00001, 0.0001, 0.0005, 0.001, 0.005
LoRA Rank 8, 16, 32, 64, 128
Search Space
It comprises three hyperparameters: model, learning rate and LoRA rank (Hu et al., 2022). We consider consider five models: 1) GPT2, 124M parameters; (Radford et al., 2019); 2) Bert-Large, 336M ; (Devlin et al., 2018); 3) Bart-Large 400 M, parameters  (Lewis et al., 2019); 4) Albert-Large, 17M parameters  (Lan et al., 2019); and 5) T5-Large, 770 M parameters  (Raffel et al., 2020). For the other two hyperparameters we also consider five different values as specified in Table 6.
Datasets
The metadataset contains predictions of models finetuned in five metadatasets for text classification: 1) IMDB (Maas et al., 2011); 2) Tweet (Maggie, 2020), 3) News (Zhang et al., 2015), 4) DBpedia (Zhang et al., 2015), 5) SST2 (Socher et al., 2013) and 6) SetFit (Tunstall et al., 2021). For every dataset, we create two versions. The first one is trained with the complete training data, while the second version is only with a subset of 10 % of the samples. All the datasets are for text classification beween 2 to 14 classes, including diverse domains such as movies, reviews, news, tweets, and text entailment data. We provide further information about the datasets in Table 7.
Figure 7: Mean error across datasets.
Metadataset Creation and Composition
We created the dataset by finetuning every model to the train split and, subsequently, saving their predictions on the validation and test split. The validation split corresponds to 20 % of the available train data. As for some datasets (SST-2 and SetFit) the test data is not completely provided by the creators, or it has hidden labels, we obtain it by using 20 % of the remaining test data. Specifically The models are fitted up to 5 epochs, using 1 GPU A100 with batch size equal to 2 and no LoRA dropout. We vary only the model, learning rate, and LoRA rank, while keeping the default hyperparameters in the Trainer object from the Transformers Library (version 4.41.0). Although we evaluate the models in a grid, some runs yielded out-of-memory errors for some configurations. In total, the metadataset contains 1134 evaluated configurations, representing around 3800 GPU hours. Additionally, we report information about the metadataset in Table 7.
Table 7: Metadataset Information
Dataset # Classes # Train Samples # Val. Samples # Test Samples # Confs (100%) # Confs. (10%)
IMDB (Maas et al., 2011) 2 20,000 5,000 25,000 125 125
Tweet (Maggie, 2020) 3 27,485 5,497 3,534 100 100
News (Zhang et al., 2015) 4 96,000 24,000 7,600 99 120
DBpedia (Zhang et al., 2015) 14 448,000 112,000 70,000 25 65
SST-2  (Socher et al., 2013) 2 43,103 13,470 10,776 125 125
SetFit  (Tunstall et al., 2021) 3 393,116 78,541 62,833 25 100
Hyperparameter Importance
We explore the importance of two hyperparameters, learning rate, and LoRA rank, by plotting the mean error as a heatmap in Figure 7. The error corresponds to the average across different models and datasets. The learning rate is an important hyperparameter, while increasing the LoRA rank does not affect the performance significantly in low learning rates. This behaviour is interesting, as it showcases that a small rank is enough for a successful finetuning in this context.
D.2 Scikit learn Pipelines
Search Space
Our primary motivation is to investigate the ensembling of automated machine learning pipelines to enhance performance across various classification tasks. To effectively study ensembling methods and benchmark different strategies, we require a diverse set of pipelines. Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al., 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers. The pipelines are structured in three stages—preprocessor, feature selector, and classifier—which allows us to systematically explore numerous configurations. This extensive and diverse search space enables us to examine the impact of ensembling on a variety of models and serves as a robust benchmark for evaluating different ensembling techniques. Detailed descriptions of the components and their hyperparameters are provided in Tables 8, 9, and 10.
Table 8: Classifiers and their hyperparameters used in the TPOT search space.
Classifier Hyperparameters
sklearn.naive_bayes.GaussianNB None
sklearn.naive_bayes.BernoulliNB
alpha (float, [1e-3, 100.0], default=50.0)
fit_prior (categorical, {True, False})
sklearn.naive_bayes.MultinomialNB
alpha (float, [1e-3, 100.0], default=50.0)
fit_prior (categorical, {True, False})
sklearn.tree.DecisionTreeClassifier
criterion (categorical, {’gini’, ’entropy’})
max_depth (int, [1, 10], default=5)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
sklearn.ensemble.ExtraTreesClassifier
n_estimators (constant, 100)
criterion (categorical, {’gini’, ’entropy’})
max_features (float, [0.05, 1.0], default=0.525)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
bootstrap (categorical, {True, False})
sklearn.ensemble.RandomForestClassifier
n_estimators (constant, 100)
criterion (categorical, {’gini’, ’entropy’})
max_features (float, [0.05, 1.0], default=0.525)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
bootstrap (categorical, {True, False})
sklearn.ensemble.GradientBoostingClassifier
n_estimators (constant, 100)
learning_rate (float, [1e-3, 1.0], default=0.5)
max_depth (int, [1, 10], default=5)
min_samples_split (int, [2, 20], default=11)
min_samples_leaf (int, [1, 20], default=11)
subsample (float, [0.05, 1.0], default=0.525)
max_features (float, [0.05, 1.0], default=0.525)
sklearn.neighbors.KNeighborsClassifier
n_neighbors (int, [1, 100], default=50)
weights (categorical, {’uniform’, ’distance’})
p (categorical, {1, 2})
sklearn.linear_model.LogisticRegression
penalty (categorical, {’l1’, ’l2’})
C (float, [1e-4, 25.0], default=12.525)
dual (categorical, {True, False})
solver (constant, ’liblinear’)
xgboost.XGBClassifier
n_estimators (constant, 100)
max_depth (int, [1, 10], default=5)
learning_rate (float, [1e-3, 1.0], default=0.5)
subsample (float, [0.05, 1.0], default=0.525)
min_child_weight (int, [1, 20], default=11)
n_jobs (constant, 1)
verbosity (constant, 0)
sklearn.linear_model.SGDClassifier
loss (categorical, {’log_loss’, ’modified_huber’})
penalty (categorical, {’elasticnet’})
alpha (float, [0.0, 0.01], default=0.005)
learning_rate (categorical, {’invscaling’, ’constant’})
fit_intercept (categorical, {True, False})
l1_ratio (float, [0.0, 1.0], default=0.5)
eta0 (float, [0.01, 1.0], default=0.505)
power_t (float, [0.0, 100.0], default=50.0)
sklearn.neural_network.MLPClassifier
alpha (float, [1e-4, 0.1], default=0.05)
learning_rate_init (float, [0.0, 1.0], default=0.5)
Table 9: Preprocessors and their hyperparameters used in the TPOT search space.
Preprocessor Hyperparameters
None None
sklearn.preprocessing.Binarizer
threshold (float, [0.0, 1.0], default=0.5)
sklearn.decomposition.FastICA
tol (float, [0.0, 1.0], default=0.0)
sklearn.cluster.FeatureAgglomeration
linkage (categorical, {’ward’, ’complete’, ’average’})
metric (categorical, {’euclidean’, ’l1’, ’l2’, ’manhattan’, ’cosine’})
sklearn.preprocessing.MaxAbsScaler None
sklearn.preprocessing.MinMaxScaler None
sklearn.preprocessing.Normalizer
norm (categorical, {’l1’, ’l2’, ’max’})
sklearn.kernel_approximation.Nystroem
kernel (categorical, {’rbf’, ’cosine’, ’chi2’, ’laplacian’, ’polynomial’,
      ’poly’, ’linear’, ’additive_chi2’, ’sigmoid’})
gamma (float, [0.0, 1.0], default=0.5)
n_components (int, [1, 10], default=5)
sklearn.decomposition.PCA
svd_solver (categorical, {’randomized’})
iterated_power (int, [1, 10], default=5)
sklearn.preprocessing.PolynomialFeatures
degree (constant, 2)
include_bias (categorical, {False})
interaction_only (categorical, {False})
sklearn.kernel_approximation.RBFSampler
gamma (float, [0.0, 1.0], default=0.5)
sklearn.preprocessing.RobustScaler None
sklearn.preprocessing.StandardScaler None
tpot.builtins.ZeroCount None
tpot.builtins.OneHotEncoder
minimum_fraction (float, [0.05, 0.25], default=0.15)
sparse (categorical, {False})
threshold (constant, 10)
Table 10: Feature selectors and their hyperparameters used in the TPOT search space.
Selector Hyperparameters
None None
sklearn.feature_selection.SelectFwe
alpha (float, [0.0, 0.05], default=0.025)
sklearn.feature_selection.SelectPercentile
percentile (int, [1, 100], default=50)
sklearn.feature_selection.VarianceThreshold
threshold (float, [0.0001, 0.2], default=0.1)
sklearn.feature_selection.RFE
step (float, [0.05, 1.0], default=0.525)
estimator (categorical, {’sklearn.ensemble.ExtraTreesClassifier’})
Estimator Hyperparameters:
   n_estimators (constant, 100)
   criterion (categorical, {’gini’, ’entropy’})
   max_features (float, [0.05, 1.0], default=0.525)
sklearn.feature_selection.SelectFromModel
threshold (float, [0.0, 1.0], default=0.5)
estimator (categorical, {’sklearn.ensemble.ExtraTreesClassifier’})
Estimator Hyperparameters:
   n_estimators (constant, 100)
   criterion (categorical, {’gini’, ’entropy’})
   max_features (float, [0.05, 1.0], default=0.525)
Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al., 2019) as the foundation for our meta-dataset. OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains. This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
Metadataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space. Each pipeline execution was constrained to a maximum runtime of $15$ minutes. During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints. For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features. We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset. This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
Appendix E Additional Results
In this Section we report additional results from our experiments:
  • Average Ranking of baselines for the Negative Log-likelihood (Table 11) and Classification Errors 12.
  • Average NLL in Scikit-learn pipelines metadataset 8.
  • Average NLL using a subset of base models selected via DivBO 13, random 14 or both 15.
Table 11: Average Ranked NLL
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
Single-Best 14.0000±0.8944 14.0000±1.0000 12.8333±1.0408 6.8833±2.3219 7.3667±2.5049 7.5361±3.8531 7.7647±3.9096
Random 16.0000±0.0000 7.6667±4.0415 9.3333±0.5774 15.2500±0.8068 15.2333±0.7739 12.0000±3.9962 13.6176±1.8669
Top5 11.3333±2.5033 9.6667±1.1547 7.0000±1.0000 6.9000±1.9360 8.3000±1.9678 6.2470±3.3951 7.2353±4.2907
Top50 10.8333±1.7224 4.6667±1.1547 4.3333±0.5774 11.4000±1.2959 11.5333±1.5025 6.8554±3.4006 6.8824±3.7397
Quick 6.1667±1.6021 6.3333±3.0551 3.6667±2.5166 4.7333±2.0331 6.0333±2.5661 6.2048±2.8531 5.9412±3.2107
Greedy 3.5000±0.8367 4.0000±2.6458 12.8333±1.0408 3.5167±1.7786 4.5667±2.4238 5.9036±3.0268 6.5882±3.2607
CMAES 12.1667±4.4460 14.0000±1.0000 12.8333±1.0408 12.9667±2.6811 13.9667±1.3322 9.9277±3.8880 6.2353±2.4630
Random Forest 7.0000±2.7568 13.0000±2.6458 14.0000±2.6458 11.9833±1.2964 9.3000±3.8699 11.4217±5.1493 8.2353±4.8157
Gradient Boosting 3.1667±4.8339 14.6667±1.5275 13.5000±2.1794 9.7000±5.2203 10.1167±5.1890 11.4217±4.9463 8.3529±4.7558
SVM 10.6667±1.3663 10.0000±7.0000 11.3333±7.2342 13.9167±1.2532 11.6833±5.0198 10.8434±4.2641 11.7941±5.5707
Linear 7.6667±2.0656 11.0000±1.0000 10.0000±2.0000 6.8000±2.8816 6.5833±3.3479 10.3855±5.2587 14.3529±3.4989
MA 11.1667±2.8577 6.6667±1.1547 4.3333±2.3094 13.2167±1.0059 12.8167±1.6000 9.0241±4.0567 8.2941±5.1813
DivBO 7.1667±3.5449 9.0000±4.5826 5.3333±5.7735 4.9167±2.5123 5.1000±2.7114 6.6145±2.9377 6.7647±2.9054
EO 5.8333±3.7103 5.3333±1.5275 7.3333±1.5275 5.8833±2.2194 5.5000±2.5052 6.2169±2.6689 7.0000±2.5739
NE-Stack (Ours) 6.0000±4.7329 1.0000±0.0000 3.3333±4.0415 3.2000±3.4381 2.3333±2.0899 10.0241±5.8351 12.2941±3.5314
NE-MA (Ours) 3.3333±2.5033 5.0000±3.0000 4.0000±1.7321 4.7333±2.0331 5.5667±1.9945 5.3735±3.2674 4.6471±2.3702
Table 12: Average Ranked Error
FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
Single-Best 12.1667±2.5626 15.0000±0.8660 13.8333±0.7638 9.8167±3.7908 9.1167±3.5591 8.9157±4.1480 9.3086±4.1030
Random 15.8333±0.4082 13.0000±0.0000 12.0000±1.0000 15.9667±0.1826 15.9667±0.1826 11.3253±5.0723 12.2716±4.3747
Top5 6.0833±4.2710 11.6667±2.0817 8.3333±2.0817 4.4000±2.8780 4.6167±2.9204 8.7048±3.6325 7.2099±4.0519
Top50 13.5000±2.2583 3.3333±1.5275 2.0000±1.0000 8.3000±4.1618 9.7333±2.6546 7.9880±4.1548 7.2222±3.6929
Quick 6.3333±4.1433 7.6667±1.1547 6.3333±2.3094 5.5833±3.5186 3.5500±2.3575 8.3253±3.1771 6.8148±3.8033
Greedy 5.0833±3.8264 11.3333±3.0551 13.8333±0.7638 9.0000±3.5012 6.9833±3.1961 8.0843±3.3467 6.9321±3.2984
CMAES 4.5833±2.2004 5.0000±3.6056 6.3333±2.0817 8.5833±3.1103 6.3167±2.1794 8.8313±3.7776 9.6605±4.5537
Random Forest 6.6667±4.4121 8.3333±4.6188 9.5000±4.4441 11.7000±3.6237 12.5000±1.9343 8.8313±4.0817 8.7778±4.5600
Gradient Boosting 10.5000±3.2711 15.6667±0.2887 14.5000±1.5000 8.4667±4.2729 12.4000±3.1139 9.1024±4.9840 9.0123±5.0812
SVM 3.5000±3.3317 8.0000±1.7321 8.6667±5.1316 10.1167±4.8382 13.3667±2.6682 7.0602±3.9296 10.5185±5.4906
Linear 7.8333±2.6394 11.6667±3.5119 11.3333±2.0817 5.7500±3.2450 7.5167±2.9870 7.7892±5.2179 9.0185±5.1705
MA 12.8333±2.4014 4.6667±0.5774 2.3333±1.5275 10.0667±3.5349 11.3667±2.9271 8.4277±4.2949 8.7531±3.7200
DivBO 9.4167±3.3229 10.6667±1.1547 12.6667±1.5275 10.0833±2.9916 7.3833±3.2262 8.1506±3.6599 8.4938±3.4228
EO 10.0000±4.6476 4.6667±3.2146 4.0000±2.0000 9.3667±2.8555 8.5000±2.6425 8.1988±4.2677 7.6852±2.8410
NE-Stack (Ours) 4.6667±3.0768 3.3333±1.1547 7.3333±0.5774 3.4500±2.3095 2.7167±2.2232 8.9639±5.4352 7.6173±5.5437
NE-MA (Ours) 7.0000±1.7889 2.0000±1.0000 3.0000±2.0000 5.3500±3.7374 3.9667±2.2967 7.3012±3.9900 6.7037±3.7855
Table 13: Average NLL for Subset of Base Models with DivBO
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
Table 14: Average NLL for Subset of Base Models randomly
Seletor FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
Table 15: Average NLL for Subset of Base Models (more baselines)
Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
Single DivBO 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9767±0.0404 0.8220±0.3507 0.8814±0.2062 0.9748±0.1415 1.0287±0.1365
Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9805±0.0339 0.7916±0.3245 0.8569±0.2178 0.9672±0.1381 1.0232±0.1550
Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9687±0.0542 0.8048±0.3271 0.8672±0.2195 0.9652±0.1353 1.0378±0.1959
Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9805±0.0339 0.7910±0.3209 0.8572±0.2174 0.9673±0.1384 1.0256±0.1601
CMAES DivBO 0.8915±0.1759 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 0.9671±0.1179 1.0250±0.1595
Random Forest DivBO 0.7932±0.1194 0.9338±0.3436 1.1678±0.2906 1.6807±1.1160 1.5507±0.8449 1.3195±0.6428 1.0534±0.1277
Gradient Boosting DivBO 0.7908±0.1848 1.4011±0.5359 1.0000±0.0000 2.4553±1.1191 2.7220±1.3007 1.2106±0.5937 1.1246±0.4052
Linear DivBO 0.7433±0.0870 0.6471±0.2272 1.0248±0.0430 0.9690±0.4209 1.1132±0.4525 1.0814±1.0056 1.0316±0.1403
SVM DivBO 0.8312±0.0943 0.7406±0.2612 1.0786±0.1362 5.2463±3.4452 4.8194±3.5169 1.5156±1.5597 2.7777±2.9458
MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9719±0.0486 0.7603±0.2876 0.8802±0.1741 0.9621±0.1306 1.0222±0.1570
Single Random 1.0067±0.0164 1.0000±0.0000 1.0540±0.0936 0.8958±0.3297 0.9786±0.2444 0.9902±0.1529 1.1314±0.4475
Top25 Random 0.8397±0.1000 0.5848±0.1980 0.7283±0.1534 2.7131±1.5510 2.5961±1.9153 1.2873±1.5773 1.0588±0.2762
Quick Random 0.7305±0.0764 0.5958±0.1917 0.7467±0.1420 0.9847±0.1836 0.9737±0.1340 1.0043±0.2497 1.0123±0.1529
Greedy Random 0.7024±0.0720 0.5783±0.1857 0.7478±0.1352 0.9171±0.0932 0.8960±0.0795 0.9762±0.1082 1.0002±0.1038
CMAES Random 1.2164±0.3851 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.1544±0.9890 0.9985±0.0954
RF Random 0.7505±0.0915 0.8325±0.2255 0.9850±0.0175 2.9838±1.7970 2.6001±1.8694 1.2006±0.4367 1.0058±0.0898
GBT Random 0.7235±0.1605 1.8377±1.4510 1.2177±0.3771 1.9521±1.2076 2.8683±1.9763 1.2329±0.6934 1.0470±0.1913
Linear Random 0.7541±0.0897 0.7400±0.2827 0.9103±0.1384 1.1776±0.3187 1.5498±0.8446 1.0417±1.0146 1.0408±0.1296
SVM Random 0.8010±0.0903 0.7767±0.3006 0.9821±0.3676 5.1993±3.4269 5.1080±3.2902 1.4479±1.4990 2.8328±3.0214
MA Random 0.9069±0.1812 0.8677±0.2292 0.7543±0.1312 4.1694±2.5844 3.0315±2.4578 1.5059±1.9654 1.4185±1.6833
NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6882±0.1602 0.6925±0.4510 0.6854±0.2801 0.9387±0.4628 4.2065±3.4678
NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.7535±0.1164 0.9291±0.0999 0.9191±0.0624 0.9922±0.2525 1.0259±0.2002
NE-Stack DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.0442±0.8691 0.7482±0.2254 0.9671±0.3746 6.9574±3.4368
NE-MA DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 0.7632±0.3319 0.8349±0.2075 0.9677±0.1455 1.0057±0.1121
Figure 8: Average Normalized NLL in Scikit-learn pipelines (Validation).

### v2

@@
-Dynamic Post-Hoc Neural Ensemblers
+Regularized Neural Ensemblers
 Abstract
 Ensemble methods are known for enhancing the accuracy and robustness of machine learning models by combining multiple base learners.
-However, standard approaches like greedy or random ensembles often fall short, as they assume a constant weight across samples for the ensemble members.
+However, standard approaches like greedy or random ensembling often fall short, as they assume a constant weight across samples for the ensemble members.
 This can limit expressiveness and hinder performance when aggregating the ensemble predictions.
-In this study, we explore employing neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively.
-Motivated by the risk of learning low-diversity ensembles, we propose regularizing the model by randomly dropping base model predictions during the training.
-We demonstrate this approach lower bounds the diversity within the ensemble, reducing overfitting and improving generalization capabilities.
-Our experiments showcase that the dynamic neural ensemblers yield competitive results compared to strong baselines in computer vision, natural language processing, and tabular data.
+In this study, we explore employing regularized neural networks as ensemble methods, emphasizing the significance of dynamic ensembling to leverage diverse model predictions adaptively.
+Motivated by the risk of learning low-diversity ensembles, we propose regularizing the ensembling model by randomly dropping base model predictions during the training.
+We demonstrate this approach provides lower bounds for the diversity within the ensemble, reducing overfitting and improving generalization capabilities.
+Our experiments showcase that the regularized neural ensemblers yield competitive results compared to strong baselines across several modalities such as computer vision, natural language processing, and tabular data.
 1  Introduction
-Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced predictive performance over single-model predictions.
-Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al., 2017).
+Figure 1: Wrong Models Per Samples Across Meta-Datasets.
+Every dark cell represents data instances where a model’s prediction is wrong.
+Different models fail on different instances, therefore, only instance-specific dynamic ensembles are optimal.
+Ensembling machine learning models is a well-established practice among practitioners and researchers, primarily due to its enhanced generalization performance over single-model predictions (Ganaie et al.,, 2022; Erickson et al.,, 2020; Feurer et al.,, 2015; Wang et al.,, 2020).
+Ensembles are favored for their superior accuracy and ability to provide calibrated uncertainty estimates and increased robustness against covariate shifts (Lakshminarayanan et al.,, 2017).
 Combined with their relative simplicity, these properties make ensembling the method of choice for many applications, such as medical imaging and autonomous driving, where reliability is paramount.
-Despite these advantages, the process of selecting post-hoc models that are both accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows.
-Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al., 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation.
+A popular set-up consists of ensembling from a pool, after training them separately, a.k.a.
+post-hoc ensembling (Purucker and Beel, 2023b, ).
+This allows users to use models trained during hyperparameter optimization.
+Despite these advantages, selecting post-hoc models that are accurate and diverse remains a challenging combinatorial problem, especially as the pool of candidate models grows.
+Commonly used heuristics, particularly in the context of tabular data, such as greedy selection (Caruana et al.,, 2004) and various weighting schemes, attempt to optimize ensemble performance based on metrics evaluated on a held-out validation set or through cross-validation.
 However, these methods face significant limitations.
-Specifically, the selection of models to include in the ensemble and the determination of optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data.
+Specifically, choosing models to include in the ensemble and determining optimal ensembling strategies (e.g., stacking weights) are critical decisions that, if not carefully managed, can lead to overfitting on the validation data.
 Although neural networks are good candidates for generating ensembling weights, few studies rely on them as a post-hoc ensembling approach.
@@
 In this work, we introduce a novel approach to post-hoc ensembling using neural networks.
-Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al., 2008).
+Our proposed Neural Ensembler dynamically generates the weights for each base model in the ensemble on a per-instance basis, a.k.a dynamical ensemble selection (Ko et al.,, 2008; Cavalin et al.,, 2013).
 To mitigate the risk of overfitting the validation set, we introduce a regularization technique inspired by the inductive biases inherent to the ensembling task.
-Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al., 2014).
+Specifically, we propose randomly dropping base models during training, inspired by previous work on DropOut in Deep Learning (Srivastava et al.,, 2014).
+Furthermore, our method is modality-agnostic, as it only relies on the base model predictions.
+For instance, for classification, we use the class predictions of the base models as input.
 In summary, our contributions are as follows:
   1. 1.
-    We propose a simple yet effective post-hoc ensembling method based on a neural network that dynamically ensembles base models.
+    We propose a simple yet effective post-hoc ensembling method based on a regularized neural network that dynamically ensembles base models and is modality-agnostic.
   2. 2.
-    To prevent the formation of low-diversity ensembles, we introduce a regularization technique that involves randomly dropping base model predictions during training.
-We demonstrate theoretically that this lower bounds the diversity of the generated ensemble.
-Additionally, we ablate its effect through various experiments.
+    To prevent the formation of low-diversity ensembles, the regularization technique randomly drops base model predictions during training.
+We demonstrate theoretically that this lower bounds the diversity of the generated ensemble, and validate its effect empirically.
   3. 3.
     Through extensive experiments, we show that Neural Ensemblers consistently select competitive ensembles across a wide range of data modalities, including tabular data (for both classification and regression), computer vision, and natural language processing.
-To promote reproducibility, we have made our code publicly available in the following anonymous repository  [^1] https://github.com/machinelearningnuremberg/NeuralEnsemblers/.
+To promote reproducibility, we have made our code publicly available in the following anonymous repository [^1] https://github.com/machinelearningnuremberg/RegularizedNeuralEnsemblers.
 We hope that our codebase, along with the diverse set of benchmarks used in our experiments, will serve as a valuable resource for the development and evaluation of future post-hoc ensembling methods.
 2  Background and Motivation
-Post-Hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$.
-These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatination of the base models predictions.
-While the base models are estimated using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
+Post-hoc ensembling uses set of fitted base models $\{z_{1},...,z_{M}\}$ such that every model outputs predictions $z_{m}(x):\mathbb{R}^{D}\rightarrow\mathbb{R}$.
+These outputs are combined by a stacking ensembler $f(z(x);\theta):=f(z_{1}(x),...,z_{M}(x);\theta):\mathbb{R}^{M}\rightarrow\mathbb{R}$, where $z(x)=[z_{1}(x),...,z_{M}(x)]$ is the concatenation of the base models predictions.
+While the base models are fitted using a training set $\mathcal{D}_{\mathrm{Train}}$, the ensembler’s parameters $\theta$ are typically obtained by minimizing a loss function on a validation set $\mathcal{D}_{\mathrm{Val}}$ such that:
 $\theta\in\operatorname*{arg\,min}_{\theta}\;\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}\mathcal{L}(f(z(x);\theta),y).$ (1)
-In the general case, this objective function can be optimized using gradient-free optimization method such as evolutionary algorithms (Purucker & Beel, 2023b) or greedy search (Caruana et al., 2004) together with a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
+In the general case, this objective function can be optimized using gradient-free optimization methods such as evolutionary algorithms (Purucker and Beel, 2023b, ) or greedy search (Caruana et al.,, 2004).
+Commonly, $f$ is a linear combination $\theta\in\mathbb{R}^{M}$ of the model outputs:
 $f(z(x);\theta)=\sum_{m}\theta_{m}z_{m}(x).$ (2)
@@
 $p(y|x)=\sum_{i}p(y|x,m)p(m),$ (3)
-which is referred to as Bayesian Model Average, and uses $\theta_{m}=p(m)$.
-In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacker model parametrized by $\beta$.
+which is referred to as Bayesian Model Average (Raftery et al.,, 1997), and uses $\theta_{m}=p(m)$.
+In the general case, the probabilistic ensembler $p(y|x)=p(y|z_{1}(x),...,z_{M}(x),\beta)$ is a stacking model parametrized by $\beta$.
 2.1  Motivating Dynamic Ensembling
-Figure 1: Wrong Models Per Samples Across Meta-Datasets.
-Every dark cell represents data instances where a model’s prediction is wrong.
-Different models fail on different instances, therefore, only instance-specific dynamic ensembles are optimal.
-We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental metadatasets.
-In reality, Equation 3 does not specify the distribution $p(m)$.
-Generally, it is safe to assume that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we dynamically select the optimal aggregation $\theta_{m}(x)=p(m|m)$ on a per-data point basis, instead of a static $\theta$.
-To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas & Erickson, 2023)), QuickTune (Computer Vision (Arango et al., 2024)), FTC (NLP-Section D.1) and NasBench 201 (NAS for Computer Vision (Dong & Yang, 2020)).
+We motivate in this Section the need for dynamic ensembling by analyzing base models’ predictions in real data taken from our experimental meta-datasets.
+Generally, the distribution $p(m)$ can take different forms.
+Dynamic ensembling assumes that the performance associated with an ensembler $f(z_{m}(x),\theta)$ is optimal if we select the optimal aggregation $\theta_{m}(x)=p(x|m)$ on a per-data point basis, instead of a static $\theta$.
+To motivate this observation, we selected four datasets from different modalities: TabRepo (Tabular data (Salinas and Erickson,, 2023)), QuickTune (Computer Vision (Arango et al., 2024a, )), FTC (Arango et al., 2024b, ) and NasBench 201 (NAS for Computer Vision (Dong and Yang,, 2020)).
 Then, we compute the per-sample error for 100 models in 100 samples.
 We report the results in Figure 1, indicating failed predictions with dark colors.
-We observe that models make different errors across samples, demonstrating the lack of optimality for static ensembling weights.
+We observe that models make different errors across samples, indicating a lack of optimality in static ensembling weights.
 3  Neural Ensemblers
@@
 In Model Averaging mode, it generates the unnormalized weights for every model, which are normalized with SoftMax.
-We use neural networks as ensemblers by training on the base model predictions $z(x)=[z_{1}(x);...;z_{M}(x)]$ as input.
-For regression, $z(x):\mathbb{R}^{D}\rightarrow\mathbb{R}^{M}$ outputs thebase models’ point predictions given by $x\in\mathbb{R}^{D}$, while for classification $z(x;c):\mathbb{R}^{D}\rightarrow[0,1]^{M}$ returns the probabilities predicted by the base models for class $c$.
-In our discussion we consider two functional modes for the ensemblers: as network outputting weights for model averaging or as a direct stacking model that outputs the prediction.
-In stacking mode for regression, we aggregate the base model point predictions using a neural network to estimate the final prediction $\hat{y}=\phi(z;\beta)$, where $\beta$ are the network parameters.
-In the model-averaging mode, the Neural Ensembler outputs the weights $\theta_{m}(z;\beta)$ to combine the model predictions as in Equation 4.
-$\hat{y}=\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}(x).$ (4)
+[…]
+$\sum_{m}\theta_{m}({z};\beta)\cdot z_{m}$ (4)
 Regardless of the functional mode, the Neural Ensembler has a different output $\hat{y}$ for regression and classification.
 In regression, the output $\hat{y}$ is a point estimation of the mean for a normal distribution such that $p(y|x;\beta)=\mathcal{N}(\hat{y},\sigma)$.
-For classification, the input is the probabilistic prediction of the base models per class $z_{m}(x;c)=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z(x),\beta)$.
+For classification, the input is the probabilistic prediction of the base models per class $z^{(c)}_{m}=p(y=c|x,m)$, while the output is a categorical distribution $\hat{y}=p(y=c|z,\beta)$.
 We optimize $\beta$ by minimizing the negative log-likelihood over the validation dataset $\mathcal{D}_{\mathrm{Val}}$ as:
-$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta).$ (5)
-3.1 Architecture
+$\min_{\beta}\mathcal{L}(\beta;D_{\mathrm{Val}})=\min_{\beta}\sum_{(x,y)\in\mathcal{D}_{\mathrm{Val}}}-\mathrm{log}\ p(y|x;\beta)$ (5)
+3.1  An Architecture with Parameter Sharing
 We discuss the architectural implementation of the Neural Ensembler for the classification case, which we show in Figure 2.
-For the stacking mode, we use an MLP that outputs the logit $\phi(z(x;c);\beta):\mathbb{R}^{M}\rightarrow[0,1]$ of each class $c$.
-Specifically, the network receives as inputs the base models’ predictions $z(x,c)$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e.
-$\phi(z;c)$.
-The model predictions per class are fed independently, enabling sharing the network parameters $\beta$.
-Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z;c)}}{\sum_{c^{\prime}}\exp^{\phi(z;c^{\prime})}}$, with $\mathrm{MLP}(z(x;c);\beta)$.
-In regression, the final prediction is the output $\phi(z(x);\beta)$.
-For model averaging mode, we use a novel architecture based on a Deep Set (Zaheer et al., 2017) embedding of the base models predictions.
-We compute the dynamic weights $\theta_{m}(z;\beta)=\frac{\exp f_{m}(z;\beta)}{\sum_{m^{\prime}}f_{m^{\prime}}(z;\beta)}$, where the unnormalized weight per model $f_{m}(z;\beta)$ is determined via two MLPs.
-The first one $\text{MLP}_{1}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ embeds the predictions per class $z(x,c^{\prime})$ into a latent dimension of size $H$, whereas the second network $\text{MLP}_{2}:\mathbb{R}^{M}\rightarrow\mathbb{R}^{H}$ aggregates the embeddings and outputs the unnormalized weights, as shown in Equation 6.
-Notice that the Neural Ensemblers’ input dimension and number of parameters are independent of the number of classes, due to our proposed parameter-sharing mechanism.
-$f(z;m,\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z(x;c^{\prime});\beta_{1}\right);\beta_{2}\right).$ (6)
-3.2 The Risk of Diversity Collapse
-In this section, we elaborate on one risk that might arise during learning ensembles, as a base model might be highly correlated with the target in the validation set, when building ensembles.
-We dub this model a Preferred Base Model.
-Definition 1 (Preferred Base Model).
-Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$.
-$z_{p}$ is the Preferred Base Model if it has the highest sample correlation to the target, i.e.
-$\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
-Proposition 1.
-If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{p}\}$, given the conditions of Definition 1.
-Sketch of Proof.
-Given that the base models predictions are uncorrelated, their correlations $\rho_{z_{m},z_{p}}=0,\forall m\neq p$, then it holds that $-\sqrt{1-\rho_{z_{p},y}^{2}}\leq\rho_{z_{m},y}\leq\sqrt{1-\rho_{z_{p},y}^{2}},\forall z_{m}\in\mathcal{F}/\{z_{p}\}$.
-We provide more details in the Appendix A.
-∎
-On the other hand, an important aspect when building ensembles is guaranteeing diversity among the models (Wood et al., 2023; Jeffares et al., 2024).
-This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al., 2022; Purucker et al., 2023).
-A common way to measure diversity is the ambiguity (Krogh & Vedelsby, 1994), which can be derived after decomposing the loss function (Jeffares et al., 2024).
-Unfortunately, even simple ensembles risk obtaining low diversity although their base models are uncorrelated.
-As we will shortly show, this happens especially when there is a preferred base model.
-Proposition 2 (Diversity Collapse).
-If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$, i.e.
-$\lim_{\rho_{z_{p},y}\to 1}\alpha=0$.
-We give a detailed proof in the Appendix A. However, we can build the intuition after noticing that given Proposition 1, $\lim_{\rho_{z_{p},y}\to 1}\bar{z}=z_{p}$.
+For the stacking mode, we use an MLP that receives as input the base models’ predictions $z^{(c)}$ for the class $c$ and outputs the corresponding predicted logit for this class, i.e.
+$\phi(z,c;\beta)$.
+The model predictions per class are fed independently into the same network, enabling the sharing of the network parameters $\beta$ as shown in Figure 6.
+Subsequently, we compute the probability $p(y=c|x)=\frac{\exp^{\phi(z,c;\beta)}}{\sum_{c^{\prime}}\exp^{\phi(z,c^{\prime};\beta)}}$, with $\phi(x,c;\beta)=\mathrm{MLP}(z^{(c)};\beta)$.
+In regression, the final prediction is the output $\phi(z;\beta)$.
+[…]
+$f_{m}(z;\beta)=\mathrm{MLP}_{2}\left(\sum_{c^{\prime}}\mathrm{MLP}_{1}\left(z^{(c^{\prime})};\beta_{1}\right);\beta_{2}\right)$ (6)
+3.2  Ensemble Diversity
+An important aspect when building ensembles is guaranteeing diversity among the models (Wood et al.,, 2023; Jeffares et al.,, 2024).
+This has motivated some approaches to explicitly account for diversity when searching the ensemble configuration  (Shen et al.,, 2022; Purucker et al.,, 2023).
+A common way to measure diversity is the ambiguity (Krogh and Vedelsby,, 1994), which can be derived after decomposing the loss function (Jeffares et al.,, 2024).
+Measuring diversity is essential as it helps to avoid overfitting in the ensemble.
+Definition 3.1 (Ensemble Diversity via Ambiguity Measure).
+[…]
 3.3  Base Models’ DropOut
@@
 5
-6      Sample masking vector $r\in\mathbb{R}^{M},r_{i}\sim\mathrm{Ber}(\gamma)$;
+[…]
 7
-8      Mask base models predictions $z_{\mathrm{drop}}(x)=r\odot z(x)$ ;
+8      Mask base models predictions $z_{\mathrm{drop}}=r\odot z$ ;
 9
 […]
-11             Compute predictions $\hat{y}=\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}};\beta\right)$ ;
+11             Compute predictions $\phi\left(\frac{1}{\gamma}z_{\mathrm{drop}}\right)$ ;
+12
 13      else
-13
-14            Compute weights $\theta\left(\frac{1}{\gamma}z_{\mathrm{drop}};m,\beta,r\right)$ using Equation 7 ;
-15
-16            Compute predictions $\hat{y}$ using Equation 4 ;
-17
+14
+15            Compute weights $\theta_{m}\left(\frac{1}{\gamma}z_{\mathrm{drop}}\right)$ using Equation 7 ;
+16
+17            Compute predictions using Equation 4 ;
+18
 19       end if
-19      Update neural ensembles parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
+20      Update parameters $\beta$ using $\nabla\mathcal{L}(\beta;\mathcal{D}_{\mathrm{Val}})$
 21 end while
-21
-22return $\beta$;
-23
+22
+23return $\beta$;
 Algorithm 1 Training Algorithm for Neural Ensemblers with Base Models’ DropOut
@@
 Moreover, it gives additional expressivity associated with neural networks.
-However, there is also a risk of overfitting and diversity collapse.
-As we showed in Section 3.2, this happens when there is a preferred model on which the ensembler mainly relies while neglecting other base model predictions.
-Although it might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples.
-Inspired by previous work (Srivastava et al., 2014), we propose to drop some base models during training forward passes.
+However, there is also a risk of overfitting caused by a low diversity.
+Although the error might effectively decrease the validation loss (Equation 5), it does not necessarily generalize to test samples.
+Inspired by previous work (Srivastava et al.,, 2014), we propose to drop some base models during training forward passes.
 Intuitively, this forces the ensembler to rely on different base models to perform the predictions, instead of merely using the preferred base model(s).
-Formally, we mask the inputs such as $r_{m}\cdot z_{m}(x)$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate.
+Formally, we mask the inputs such as $r_{m}\cdot z_{m}$, where $r_{m}\sim\mathrm{Ber}(\gamma)$, where $\mathrm{Ber}(\gamma)$ is the Bernoulli distribution with parameter $\gamma$ with represents the probability of keeping the base model, while $\delta=1-\gamma$ represents the DropOut rate.
 We also mask the weights when using model averaging:
-$\theta(z;m,\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
-As DropOut changes the scale of the inputs, the weight scaling rule should be applied during inference by multiplying the dropped variables by the retention probability $\gamma$.
+$\theta_{m}(z;\beta,r)=\frac{r_{m}\cdot\exp{f_{m}}({z};\beta)}{\sum_{m^{\prime}}r_{m^{\prime}}\cdot\exp{f_{m^{\prime}}}({z};\beta)}.$ (7)
+As DropOut changes the scale of the inputs, we should apply the weight scaling rule during inference by multiplying the dropped variables by the retention probability $\gamma$.
 Alternatively, we can scale the variables during training by $\frac{1}{\gamma}$.
@@
 It has two modes, acting as a direct stacker or as a model averaging ensembler.
-We demonstrate that base models’ DropOut avoids diversity collapse by upper bounding the diversity as stated in Proposition 6, even for the simplest ensembling case.
-Proposition 3 (Avoiding Diversity Collapse).
-As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
-Sketch of Proof.
-We want to compute $\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$.
-By using $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$, and assuming, without loss of generality, that the predictions are standardized, we obtain $\mathbb{V}(r\cdot z_{m})=\gamma$.
-This lead as to $\lim_{\rho_{z_{p},y}\to 1}\alpha=1-\gamma$, after following a procedure similar to Proposition 2.
-We provide the complete proof in Appendix A.
+We demonstrate that base models’ DropOut lower bounds the diversity for a simple ensembling case in Proposition A.1.
+Definition 3.2 (Preferred Base Model).
+Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$.
+$z_{p}$ is the Preferred Base Model if it has the highest sample correlation to the target, i.e.
+$\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
+Proposition 3.3 (Diversity Lower Bound).
+As the correlation of the preferred model increases $\rho_{p_{m},y}\rightarrow 1$, the diversity decreases $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
+[…]
 4  Experiments and Results
-4.1 Proof-of-Concept: Ensemble of Quadratic Functions
-To further highlight the importance of the dynamical ensemble with diverse base models, we propose a simple regression problem with a third-degree polynomial with the ground truth function $y_{\mathrm{true}}=1.33x^{3}-0.77x^{2}-0.31x-1$, where $x\in\left[-1,1\right]$.
-We fit three second-degree polynomials on different subsets of the training dataset, to obtain three base functions $z_{1}(x),z_{2}(x),z_{3}(x)$.
-As ensemblers we consider i) a static model average $\sum_{i}\theta_{m}z_{m}(x)$, ii) a dynamic model stacker $\phi(z(x);\beta)$ and iii) model average $\sum_{m}\theta_{m}(z(x);\beta)\cdot z_{m}(x)$.
-In this experiment, $\phi(\cdot)$ and $\theta_{m}(\cdot)$ are two-layer MLPs with 10 hidden neurons with parameters $\beta$ and trained on the validation data.
-Figure 3 shows the specific data, base functions, and learned ensemblers.
-The rightmost plot shows that the dynamic model averaging (MSE=$0.0101$) and stacker (MSE=$0.0055$), by expressing more complex functions, better model the ground truth than the static ensemble (MSE=$0.1662$).
-Figure 3: Intuition Example for Dynamical Ensembling.
-When considering three base models $z_{1}(x),z_{2}(x),z_{3}(x)$, the dynamical ensemblers $\phi(\cdot)$ and $\theta_{m}(\cdot)$ depending on $x$ achieve better performance than a static model average, as observable in the right-most figure.
+In this section, we provide empirical evidence of the effectiveness of our approach.
 4.1  Experimental Setup
@@
 These meta-datasets cover diverse data modalities, including Computer Vision, Tabular Data, and Natural Language Processing.
-Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that require model fitting.
+Additionally, we evaluate the method on datasets without pre-computed predictions to assess the performance of ensembling methods that rely on fitted models during the evaluation.
 Table 11 reports the main information related to these datasets.
 Particularly for Nasbench, we created 2 versions by subsampling 100 and 1000 models.
-The metadataset for Finetuning Text Classifiers (FTC) was generated by ourselves to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al., 2019), Bert (Devlin et al., 2018) and Bart (Lewis et al., 2019).
+The meta-dataset Finetuning Text Classifiers (FTC) (Arango et al., 2024b, ) contains the predictions of several language models to evaluate ensembling techniques on text classification tasks by finetuning language models such as GPT2 (Radford et al.,, 2019), Bert (Devlin et al.,, 2018) and Bart (Lewis et al.,, 2019).
 We also generate a set of fitted Scikit-Learn Pipelines on classification datasets.
 In this case, we stored the pipeline in memory, allowing us to evaluate our method in practical scenarios where the user has fitted models instead of predictions.
-We detailed information about the creation of these two metadatasets in Appendix D.
-Information about each dataset lies in the respective referred work under the column Task Information in Table 1.
-Table 1: Metadatasets Information.
-Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
-Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
-Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
-QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
-QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
-FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
-TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
-TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
-Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
+We detailed information about the creation of this meta-dataset in Appendix O.
+Altogether, the meta-datasets comprise over 240 datasets in total.
+The information about each data set lies in the referred work under the column Task Information in Table 11.
+Table 1: Average Normalized NLL across Metadatasets.
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
+Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
+Random 1.5450±0.5289 0.6591±0.2480 0.7570±0.2900 6.8911±3.1781 5.8577±3.2546 1.7225±1.9645 1.8319±2.1395
+Top5 0.8406±0.0723 0.6659±0.1726 0.6789±0.3049 1.5449±1.8358 1.1496±0.3684 1.0307±0.5732 0.9939±0.0517
+Top50 0.8250±0.1139 0.5849±0.2039 0.6487±0.3152 3.3068±2.6197 3.0618±2.2960 1.0929±1.0198 1.0327±0.2032
+Quick 0.7273±0.0765 0.5957±0.1940 0.6497±0.3030 1.1976±1.1032 0.9747±0.2082 0.9860±0.2201 1.0211±0.1405
+Greedy 0.6943±0.0732 0.5785±0.1972 0.7617±0.3435 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
+CMAES 1.2356±0.5295 1.0000±0.0000 1.0000±0.0000 4.1728±2.8724 4.6474±3.0180 1.3487±1.3390 1.0281±0.1977
+Random Forest 0.7496±0.0940 0.8961±0.3159 0.9340±0.4262 3.7033±2.8145 2.2938±2.2068 1.2655±0.4692 1.0030±0.0871
+Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2575±0.4460 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
+SVM 0.7990±0.0909 0.7744±0.2967 0.9358±0.5706 5.4377±3.3807 4.0019±3.6601 1.3884±1.4276 2.7975±3.0219
+Linear 0.7555±0.0898 0.7400±0.2827 0.8071±0.2206 1.3960±1.2334 1.1031±0.7038 1.1976±1.1024 3.1488±3.2813
+XGBoost 0.8292±0.1434 0.7389±0.2326 0.9092±0.5304 3.7822±3.1194 2.6119±2.3911 1.7697±1.4672 1.2580±0.4875
+CatBoost 0.6887±0.0953 0.8092±0.2513 0.9512±0.5083 2.6262±2.6482 2.4145±1.8989 1.2570±1.2859 1.0454±0.1550
+LightGBM 0.7973±0.1946 3.6004±2.5822 5.3943±4.7980 3.0378±2.7945 3.6860±3.2856 1.8298±1.1596 1.6250±2.1651
+Akaike 0.8526±0.1403 0.5838±0.2031 0.6485±0.3166 3.1574±2.5898 2.6888±2.0620 1.0930±1.0203 1.0221±0.1793
+MA 0.9067±0.1809 0.5970±0.2034 0.6530±0.3028 4.7921±3.0780 4.0168±2.8560 1.4724±1.9401 1.3342±1.3515
+DivBO 0.7695±0.1195 0.7307±0.3061 0.6628±0.3435 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
+EO 0.7535±0.1156 0.5801±0.2051 0.6911±0.2875 1.3702±1.6389 0.9649±0.2980 1.0979±1.0289 1.0183±0.0993
+NE-Stack 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
+NE-MA 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
 Baselines.
 We compare the Neural Ensemblers (NE) with other common and competitive ensemble approaches.
-1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al., 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker & Beel, 2023b) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3.
-We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al., 2022), and 9) Ensemble Optimization (EO) (Levesque et al., 2016).
-Finally, we report results by using common ML models as stackers, such as 10) SVM, 11) Random Forest, 12) Gradient Boosting, and 13) Logistic/Linear Regression.
-We used the default configurations provided by Sckit-learn (Pedregosa et al., 2011) for these stackers.
-The input to the models is the concatenation of all the base models’ predictions.
-We concatenated the probabilistic predictions from all the classes in the classification tasks.
-This sometimes produced a large dimensional input space.
-Finally, we include models from Dynamic Ensemble Search (DES) literature that are in the DESlib library (Cruz et al., 2020) such as KNOP (Cavalin et al., 2013), KNORAE (Ko et al., 2008) and MetaDES (Cruz et al., 2015).
+1) Single best selects the best model according to the validation metric; 2) Random chooses randomly $N=50$ models to ensemble, 3) Top-N ensembles the best $N$ models according to the validation metric; 4) Greedy creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work (Caruana et al.,, 2004); 5) Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric; 6) CMAES (Purucker and Beel, 2023b, ) uses an evolutive strategy with a post-processing method for ensembling, 7) Model Average (MA) computed the sum of the predictions with constant weights as in Equation 3.
+We also compare to methods that perform ensemble search iteratively via Bayesian Optimization such as 8) DivBO (Shen et al.,, 2022), and 9) Ensemble Optimization (EO) (Levesque et al.,, 2016).
+We also report results by using common ML models as stackers, such as SVM and Random Forest.
+Finally, we include models from Dynamic Ensemble Search (DES) literature such as KNOP (Cavalin et al.,, 2013), KNORAE (Ko et al.,, 2008) and MetaDES (Cruz et al.,, 2015).
+We provide further details on the baselines setup in Appendix E.
+Figure 3: Results on Scikit-Learn Pipelines.
 Neural Ensemblers’ Setup.
@@
 Additionally, we used the Adam optimizer and a network with four layers, $32$ neurons, and a probability of keeping base models $\gamma=0.25$, or alternatively a DropOut rate $\delta=0.75$.
-Notice that the architecture of the ensemblers slightly varies depending on the mode (Stacking or Model Average).
+Notice that the architecture of the ensemblers slightly varies depending on the mode, Stacking or Model Average (MA).
 For the Stacking mode, we use an MLP with four layers and $32$ neurons with ReLU activations.
 For MA mode, we use two MLPs as in Equation 6: 1) $\textrm{MLP}_{1}$ has $3$ layers with $32$ neurons, while 2) $\textrm{MLP}_{2}$ has one layer with the same number of neurons.
-Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune metadataset (extended version).
-4.3 Research Questions and Associated Experiments
-Table 2: Average Normalized Error.
-FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Class (AUC)
+Although changing some of these hyperparameters might improve the performance, we keep this setup constant for all the experiments, after checking that the Neural Ensemblers perform well in a subset of the Quick-Tune meta-dataset (extended version).
+Table 2: Average Normalized Error across Metadatasets.
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
 Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
@@
 Quick 0.9494±0.0371 0.6524±0.0436 0.5787±0.1510 0.7575±0.2924 0.7879±0.2623 0.9869±0.1667 0.9054±0.2232
-Greedy 0.9494±0.0374 0.7400±0.1131 1.0000±0.0000 0.9863±0.4286 0.9297±0.1435 0.9891±0.1693 0.9090±0.2197
+Greedy 0.9494±0.0374 0.7400±0.1131 0.6033±0.1572 0.9863±0.4286 0.9297±0.1435 0.9891±0.1693 0.9090±0.2197
 CMAES 0.9489±0.0392 0.6401±0.0343 0.5797±0.1575 1.0319±0.5000 0.9086±0.1121 0.9935±0.1953 1.1878±1.1457
@@
 Linear 0.9609±0.0347 0.7891±0.1978 0.7782±0.1941 0.7333±0.4457 0.9291±0.3580 0.9776±0.2844 1.0329±0.4022
+XGBoost 0.9784±0.0334 0.7886±0.0903 0.7183±0.2628 3.3875±3.1419 2.1067±1.9570 1.3310±1.4402 1.1490±1.0202
+CatBoost 1.0271±0.0952 0.9247±0.0652 0.8526±0.2627 1.1858±0.5737 1.4011±0.7817 1.1507±1.0190 1.0179±0.2395
+LightGBM 0.9774±0.0369 1.2866±0.6738 1.2278±0.8503 2.4846±2.5747 3.1337±2.9066 1.0954±0.4326 1.1228±0.9022
+Akaike 1.0242±0.0723 0.6328±0.0384 0.5667±0.1578 1.0323±0.8817 1.0346±0.4554 0.9832±0.1784 1.1146±1.0218
 MA 1.0709±0.0845 0.6381±0.0349 0.5610±0.1490 1.1548±0.8465 1.2173±0.6107 1.0917±1.0135 0.9977±0.2278
@@
 EO 1.0208±0.1159 0.6365±0.0445 0.5704±0.1619 1.0185±0.6464 1.0367±0.4394 1.0851±1.0136 0.9377±0.2310
-NE-Stack (Ours) 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
-NE-MA (Ours) 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
+NE-Stack 0.9491±0.0451 0.6331±0.0378 0.5836±0.1592 0.6104±0.3656 0.7545±0.2960 1.0440±0.3309 1.0035±0.5295
+NE-MA 0.9527±0.0402 0.6307±0.0363 0.5621±0.1483 0.8297±0.4974 0.8236±0.2240 0.9592±0.2144 0.9028±0.2157
+4.2  Research Questions and Associated Experiments
 RQ 1: Can Neural Ensemblers outperform other common and competitive ensembling methods across data modalities?
-Figure 4: Results on Scikit-Learn Pipelines.
 Experimental Protocol.
-To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the metadatasets.
+To answer this question, we compare the neural ensembles in stacking and averaging mode to the baselines across all the meta-datasets.
 We run every ensembling method three times for every dataset.
 In all the methods we use the validation data for fitting the ensemble, while we report the results on the test split.
-Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and error.
+Specifically, we report the average across datasets of two metrics: negative log-likelihood (NLL) and classification error.
 For the tabular classification, we compute the ROC-AUC.
@@
 Results.
-The results reported in Table 2 and Table 3 show that our proposed neural networks are competitive post-hoc ensemblers.
+The results reported in Table 2 and Table 1 show that our proposed regularized neural networks are competitive post-hoc ensemblers.
 In general, we observe that the Neural Ensemblers variants obtain either the best (in bold) or second best (italic) performance across almost all meta-datasets and metrics.
-Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class.
-This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al., 2020).
+Noteworthily, the greedy approach is very competitive, especially for the FTC and TR-Class meta-datasets.
+This is coherent with previous work supporting greedy ensembling as a robust method for tabular data (Erickson et al.,, 2020).
 We hypothesize that dynamic ensembling contributes partially to the strong results for the Neural Ensemblers.
@@
 To understand this, we compare to Dynamic Ensemble Selection (DES) methods.
-Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines metadataset, as we can easily access the fitted models.
+Specifically, we use KNOP, MetaDES, and KNORAE, and evaluate all methods in Scikit-learn Pipelines meta-dataset, as we can easily access the fitted models.
 We report the results of the test split in Figure 3, where we distinguish among four types of models to facilitate the reading: Neural, DES, Stacking and Constant.
 We can see that Neural Ensemblers are the most competitive approaches, especially on stacking mode.
-Additionally, we report the metrics on the validation split in Figure 8 (Appendix E), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust in this sense.
-Table 3: Average Normalized NLL.
-FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
-Single-Best 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
-Random 1.5450±0.5289 0.6591±0.2480 0.7570±0.2900 6.8911±3.1781 5.8577±3.2546 1.7225±1.9645 1.8319±2.1395
-Top5 0.8406±0.0723 0.6659±0.1726 0.6789±0.3049 1.5449±1.8358 1.1496±0.3684 1.0307±0.5732 0.9939±0.0517
-Top50 0.8250±0.1139 0.5849±0.2039 0.6487±0.3152 3.3068±2.6197 3.0618±2.2960 1.0929±1.0198 1.0327±0.2032
-Quick 0.7273±0.0765 0.5957±0.1940 0.6497±0.3030 1.1976±1.1032 0.9747±0.2082 0.9860±0.2201 1.0211±0.1405
-Greedy 0.6943±0.0732 0.5785±0.1972 1.0000±0.0000 0.9025±0.2378 0.9093±0.1017 0.9665±0.0926 1.0149±0.1140
-CMAES 1.2356±0.5295 1.0000±0.0000 1.0000±0.0000 4.1728±2.8724 4.6474±3.0180 1.3487±1.3390 1.0281±0.1977
-Random Forest 0.7496±0.0940 0.8961±0.3159 0.9340±0.4262 3.7033±2.8145 2.2938±2.2068 1.2655±0.4692 1.0030±0.0871
-Gradient Boosting 0.7159±0.1529 1.7288±1.2623 1.2764±0.4787 1.9373±1.2839 2.6193±2.3159 1.4288±1.2083 1.0498±0.2128
-SVM 0.7990±0.0909 0.7744±0.2967 0.9358±0.5706 5.4377±3.3807 4.0019±3.6601 1.3884±1.4276 2.7975±3.0219
-Linear 0.7555±0.0898 0.7400±0.2827 0.8071±0.2206 1.3960±1.2334 1.1031±0.7038 1.1976±1.1024 3.1488±3.2813
-MA 0.9067±0.1809 0.5970±0.2034 0.6530±0.3028 4.7921±3.0780 4.0168±2.8560 1.4724±1.9401 1.3342±1.3515
-DivBO 0.7695±0.1195 0.7307±0.3061 0.7125±0.3982 1.2251±1.0293 0.9430±0.2036 1.0023±0.3411 1.0247±0.1473
-EO 0.7535±0.1156 0.5801±0.2051 0.6911±0.2875 1.3702±1.6389 0.9649±0.2980 1.0979±1.0289 1.0183±0.0993
-NE-Stack (Ours) 0.7562±0.1836 0.5278±0.2127 0.6336±0.3456 0.7486±0.6831 0.6769±0.2612 1.3268±0.7498 1.2379±0.4083
-NE-MA (Ours) 0.6952±0.0730 0.5822±0.2147 0.6522±0.3131 1.0177±0.5151 0.9166±0.0936 1.0515±1.0003 0.9579±0.0777
-RQ 2: Do Neural Ensemblers need a strong group of base models, i.e. found using Bayesian Optimization?
+Additionally, we report the metrics on the validation split in Figure 6 (Appendix F), where we observe that some dynamic ensemble approaches such as Gradient Boosting (GBT), Random Forest (RF) and KNORAE exhibit overfitting, while Neural Enemblers are more robust.
+RQ 2: What is the impact of the DropOut regularization scheme?
 Experimental Protocol.
-Practitioners use some methods such as greedy ensembling as post-hoc ensemblers, i.e., they consider a set of models selected by a search algorithm such as Bayesian Optimization as base learners.
-DivBO enhances the Bayesian Optimization by accounting for the diversity in the ensemble in the acquisition function.
-We run experiments to understand whether the Neural Ensemblers’ performance depends on a strong subset of $50$ base models selected by DivBO, and whether it can help other methods.
-We conduct additional experiments by randomly selecting $50$ models to understand the impact and significance of merely using a smaller set of base models.
-We normalize the base of the metric on the single-best base model from the complete set contained in the respective dataset.
+To understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$.
+We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset.
+In this setup, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
 Results.
-We report in Table 4 the results with the two selection methods (random and DivBO) using a subset of common baselines, where we normalize using the metric of the single-best from the whole set of models.
-We limit the number of baselines for brevity, and the extended version includes more baselines in Table 15.
-We also include results in separate tables for the two selection methods (Appendix E).
-We can compare directly with the results in Table 3.
-We observe that reducing the number of base models with DivBO negatively affects the performance of the Neural Ensemblers.
-Surprisingly, randomly selecting the subset of base models improves the results in two metadatasets (TR-Class and NB-1000).
-We hypothesize that decreasing the number of base models is beneficial for these metadatasets.
-With over $1000$ base models available, the likelihood of identifying a preferred model and overfitting the validation data increases in these metadatasets.
-Naturally, decreasing the number of base models can also be detrimental for the Neural Ensemblers, as this happens for some metadatasets such as TR-Reg and QT-Micro.
-In contrast to the Neural Ensemblers, selecting a subset of strong models with DivBO improves the performance for some baselines such as Model Averaging (MA) or TopK ($K=25$).
-In other words, it works as a preprocessing method for these ensembling approaches.
-Overall, the results in Tables 4 and 15 demonstrate that Neural Ensemblers do not need a strong group of base models to achieve competitive results.
-Table 4: Average Normalized NLL with a Subset of Base Models.
-Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
-Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
-Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
-Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
-Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
-Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
-Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
-MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
-Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
-Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
-Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
-Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
-MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
-NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
-NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
-NE-Stack (Ours) Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
-NE-MA (Ours) Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
-RQ 3: What is the impact of the DropOut regularization scheme?
-Experimental Protocol.
-Firstly, we do a small experiment to observe how the DropOut rate affects the weights when using Neural Ensemblers in MA mode.
-For this, we use QT-Micro datasets, and save each model’s weights $\theta_{m}$ and error after training the neural network.
-Subsequently, to understand how much the base learners DropOut helps the Neural Ensemblers, we run an ablation by trying the following values for the DropOut rate $\delta\in\{0.0,0.1,\dots,0.9\}$.
-We compute the average NLL for three seeds per dataset and divide this value by the one obtained for $\delta=0.0$ in the same dataset.
-Therefore, we realize that a specific DropOut rate is improving over the default network without regularization if the normalized NLL is below 1.
-Results.
-We show how the mean of the weights per model is related to the mean error of the models for different DropOut rates in Figure 5.
-When there is no DropOut some weights are close to one, i.e.
-they are preferred models.
-As we increase the value, many models with high weights decrease.
-If the rate is very high (e.g.
-$0.9$), we will have many models contributing to the ensemble, with weights different from zero.
 Our ablation study demonstrates that non-existing or high DropOut are detrimental to the Neural Ensembler performance in general.
@@
 In general, we observe that Neural Ensemblers obtain better performance when using base models’ DropOut.
-Figure 5: Mean weights assigned to the base models decrease with DropOut rate. Every data point is a model. The errors and weights are the mean values across many datasets for every model.
+Furthermore, we show how the mean weight per model is related to the mean error of the models for different DropOut rates in Figure 15 in the Appendix N.
 Figure 4: Ablation of the DropOut rate.
+RQ 3: How sensible are the Neural Ensemblers to its hyperparameters?
+Experimental Protocol.
+While our proposed Neural Ensembler (NE) uses MLPs with 4 layers and 32 neurons per layer by default, it is important to understand how sensitive the NE’s performance is to changes in its hyperparameters.
+To this end, we conduct an extensive ablation study by varying the number of layers $L\in\{1,2,3,4,5\}$ and the number of neurons per layer (hidden dimension) $H\in\{8,16,32,64,128\}$ in the stacking mode of the NE.
+For each configuration, we compute the Negative Log Likelihood (NLL) on every task in the metadatasets and normalize these values by the performance obtained with the default hyperparameters ($L=4$, $H=32$).
+This normalization allows us to compare performance changes across different tasks and metadatasets, accounting for differences in metric scales.
+A normalized NLL value below one indicates improved performance compared to the default setting.
+Results.
+Our findings indicate that there is no single hyperparameter configuration that is optimal across all datasets.
+However, our default configuration ($L=4$, $H=32$) strikes a balance, providing robust performance across diverse tasks without the need for extensive hyperparameter tuning, and can be effectively applied in various settings without the need for dataset-specific hyperparameter optimization.
+Significance on Datasets with a lot of Classes
+Given the high performance between Greedy and NE-MA, we wanted to understand when the second one would obtain strong significant results.
+We found that NE-MA is particularly well-performing in datasets with a large number of classes.
+Given Table 11, we can see that four meta-datasets have a high ($>10$) number of classes.
+We selected these metadatasets (NB(100), NB(1000), QT-Micro, QT-Mini), and plotted the significance compared to Greedy and Random Search.
+The results reported in Figure 9 demonstrate that our approach is significantly better than Greedy in these metadatasets.
+Additional research questions
+We further discuss interesting research questions in the Appendix due to the limited space.
+Some of these include: i) how much the validation data size affects the Neural Ensembler (Appendix J), ii) what happens if we train the Neural Ensemblers on a merged split containing both training and validation data (Appendix K), and iii) the disadvantages of Neural Ensemblers acting on the original input space (Appendix L).
+Figure 5: Ablation study of Neural Ensembler hyperparameters: number of neurons per layers (a) and number of layers (b).
+Normalized NLL values below one indicate improved performance over the default setting ($L=4$, $H=32$).
 5  Related Work
 Ensembles for Tabular Data.
-For tabular data, ensembles are known to perform better than individual models (Sagi & Rokach, 2018; Salinas & Erickson, 2023).Therefore, ensembles are often used in real-world applications (Dong et al., 2020), to win competitions (Koren, 2009; Kaggle, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker & Beel, 2023b; Purucker et al., 2023).
-Methods like Bagging (Breiman, 1996) or Boosting (Freund et al., 1996) are often used to boost the performance of individual models.
-In contrast, post-hoc ensembling (Shen et al., 2022; Purucker & Beel, 2023a) aggregates the predictions of an arbitrary set of fitted base models.
-Post-hoc ensembles are build by stacking (Wolpert, 1992; Van der Laan et al., 2007), ensemble selection (a.k.a.
-pruning) (Caruana et al., 2004; Tsoumakas et al., 2009), dynamic ensemble selection (Ko et al., 2008; Britto Jr et al., 2014), or through a systematic search for an optimal ensemble (Levesque et al., 2016; Shen et al., 2022; Poduval et al., 2024).
+For tabular data, ensembles are known to perform better than individual models (Sagi and Rokach,, 2018; Salinas and Erickson,, 2023).Therefore, ensembles are often used in real-world applications (Dong et al.,, 2020), to win competitions (Koren,, 2009; Kaggle,, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker and Beel, 2023b, ; Purucker et al.,, 2023).
+Methods like Bagging (Breiman,, 1996) or Boosting (Freund et al.,, 1996) are often used to boost the performance of individual models.
+In contrast, post-hoc ensembling (Shen et al.,, 2022; Purucker and Beel, 2023a, ) aggregates the predictions of an arbitrary set of fitted base models.
+Post-hoc ensembles can be built by stacking (Wolpert,, 1992) ensemble selection (a.k.a.
+pruning) (Caruana et al.,, 2004) , or through a systematic search for an optimal ensemble (Levesque et al.,, 2016).
 Ensembles for Deep Learning.
-Ensembles of neural networks (Hansen & Salamon, 1990; Krogh & Vedelsby, 1994; Dietterich, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation.
-Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al., 2017) being the most popular one, which involve independently training multiple initializations of the same network.
-Extensive empirical studies (Ovadia et al., 2019; Gustafsson et al., 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al., 2015; Gal & Ghahramani, 2016; Welling & Teh, 2011).
-Mixture-of-Experts.
-Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017), where one network is trained with specialized sub-modules that are activated based on the input data.
-Batch ensembles (Wen et al., 2020; Wenzel et al., 2020) are also closely related, as they aim to capture diverse model behaviors by simultaneously training multiple models with shared weights but different random projections.
-In contrast to MoE and batch ensembles, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches.
-Ours additionally considers a regularization strategy that leads to provably more diverse ensembles.
+Ensembles of neural networks (Hansen and Salamon,, 1990; Krogh and Vedelsby,, 1994; Dietterich,, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation.
+Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al.,, 2017) being the most popular one, which involves independently training multiple initializations of the same network.
+Extensive empirical studies (Ovadia et al.,, 2019; Gustafsson et al.,, 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al.,, 2015; Gal and Ghahramani,, 2016; Welling and Teh,, 2011).
 Dynamic Ensemble Selection.
 Our Neural Ensembler is highly related to dynamic ensemble selection.
-Both dynamically aggregate the predictions of base models per instance (Cavalin et al., 2013; Ko et al., 2008).
-Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al., 2018; 2020).
+Both dynamically aggregate the predictions of base models per instance (Cavalin et al.,, 2013; Ko et al.,, 2008).
+Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al.,, 2018, 2020).
 In contrast, we use an end-to-end trained neural network to select and weight the base models per instance.
+Mixture-of-Experts.
+Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al.,, 1991; Jordan and Jacobs,, 1993; Shazeer et al.,, 2017), where one network is trained with specialized sub-modules that are activated based on the input data.
+Alternatively, we could include a layer, aggregating predictions by encouraging diversity (Zhang et al.,, 2020).
+In contrast to these approaches, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches.
 6  Conclusions
@@
 Our empirical results demonstrated that Neural Ensemblers consistently form competitive ensembles across diverse data modalities, including tabular data (classification and regression), computer vision, and natural language processing.
-In the future, we aim to explore in-context learning (Brown et al., 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input.
-We discuss broader impact and limitations in Appendix B.
-Reproducibility Statement
-All code and datasets used in this paper are publicly available at https://github.com/machinelearningnuremberg/NeuralEnsemblers/.
-We provide the code for running all the experiments, including baselines on all meta-datasets.
-We also provide instructions for the environment setup to ensure reproducibility.
-We will deanonymize the link upon acceptance.
-We further provide all proofs in Appendix A.
-Moreover, we ensured reproducibility by testing our approach on many meta-datasets and across several domains, for which details can be found in Appendix D.
-We also compared our method to many baselines commonly used in ensembling, described in Section 4.3.
-Lastly, we detail additional results and extended data tables in Appendix E.
 Acknowledgments
@@
 Appendix
+We want to summarize here all the appendix sections:
+  • Section A presents the proofs of propositions in the main paper.
+  • Section B discusses the limitations of our proposed method and its broader impacts.
+  • Section D further details research similar to our work.
+  • Section O explains the process of gathering and preparing the FTC and Scikit-learn Pipelines metadatasets used in our experiments.
+  • Section F includes supplementary tables and figures, such as average ranks and detailed performance metrics, that support and expand upon the main experimental results reported in the paper.
+  • Section H includes he computational cost associated with our method compared to baseline approaches, including runtime evaluations and discussions on efficiency.
+  • Section I includes the results of a proof-of-concept experiment using overparameterized base models (e.g., 10th-degree polynomials), demonstrating the effectiveness of our Neural Ensembler even when base models have high capacity.
+  • Section G includes the Critical Difference diagrams corresponding to our main results, illustrating the statistical significance of performance differences among methods and how to interpret these diagrams.
+  • Section J explores the impact of using different amounts of validation data to train the Neural Ensembler, assessing its sample efficiency and how performance scales with varying data sizes.
+  • Section K explores the effect of merging the training and validation datasets on the performance of both base models and ensemblers.
+  • Section L explores an alternative formulation of the Neural Ensembler that operates on the original input space rather than on the predictions of the base model, including experimental results and discussions of its effectiveness.
+  • In section M, we study whether the performance of the models changes when using base models found by Bayesian Optimization or randomly.
 Appendix A Proofs
-Definition 2 (Preferred Base Model).
-Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$.
-$z_{p}$ is the Preferred Base Model if it has the highest sample correlation with respect the target, i.e.
-$\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
-Proposition 4.
-If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{k}\}$, given the conditions of Definition 1.
-Proof.
-Let $\rho_{z_{p},z_{m}},\rho_{z_{m},y},\rho_{z_{p},y}$ be the correlations among the preferred model $z_{p}$, a non-preferred model $z_{m}\neq z_{m}$ and the target $y$, then the correlation matrix is given by,
-$C=\begin{pmatrix}1&\rho_{z_{p},z_{m}}&\rho_{z_{p},y}\\ \rho_{z_{p},z_{m}}&1&\rho_{z_{m},y}\\ \rho_{z_{p},y}&\rho_{z_{m},y}&1\\ \end{pmatrix}$ (8)
-The determinant of the correlation matrix is given by,
-$\mathrm{Det}(C)$ $=1-\rho_{z_{p},z_{m}}^{2}-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}+2\rho_{z_{p},z_{m}}\rho_{z_{p},y}\rho_{z_{m},y}$ (9)
-$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}$ (10)
-$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}\geq 0$ (11)
-$\rho_{z_{m},y}$ $\leq\sqrt{1-\rho_{z_{p},y}^{2}},$ (12)
-where we used the fact that $\rho_{z_{p},z_{m}}=0$, in Equation 10 and that the correlation matrix should be semidefinite in Equation 11, thus its determinant should hold $\mathrm{Det}(C)\geq 0$.
-This means that $\epsilon=\sqrt{1-\rho_{z_{p},y}^{2}}$ is an upper bound for the absolute value of the correlation $\rho_{z_{m},y}$.
-The limit of the upper bound $\epsilon$ as $\rho_{z_{p},y}\to 1$ is 0 and thus $\rho_{z_{m},y}=0$,
-$\lim_{\rho_{z_{p},y}\to 1}$ $\epsilon=\lim_{\rho_{z_{p},y}\to 1}\sqrt{1-\rho_{z_{p},y}^{2}}=0$ (13)
-$\lim_{\rho_{z_{p},y}\to 1}$ $\rho_{z_{m},y}=0.$ (14)
-Since $\rho_{z_{m},y}$ is bounded by $-\epsilon\leq\rho_{z_{m},y}\leq\epsilon$.
-Proposition 5 (Diversity Collapse).
-If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$.
-Proof.
-Firstly, we show that under conditions of Definition 1, $\theta_{m}=\hat{\rho}_{z_{m},y}$, where $\hat{\rho}_{z_{m},y}$ is the sample correlation.
-Assume that we have a matrix of evaluations $X\in\mathrm{R}^{N\times M}$, with components $x_{n,m}$ indicating the prediction of the $m$-th model for the $n$-th instance.
-If $Y\in\mathbb{R}^{N\times 1}$ is the ground-truth, then the ensemble weights $\Theta\in\mathbb{R}^{M\times 1}$ can be computed with closed form solution of the objective $\min_{\Theta}||Y-X\Theta||^{2}$ as: $\Theta=(X^{T}X)^{-1}X^{T}Y$.
-Without loss of generality, we assume that the random variables are standardized.
-Given that every row of $X$ have samples from uncorrelated random variables, then $X^{T}X\approx I$, denoting $I$ as the identity matrix, thus
-$\Theta=X^{T}Y$ (15)
-Every component of $\Theta$ can expressed as summatorias $\theta_{m}=\sum_{n}x_{n,m}y_{n}$.
-Note that $x_{n,m}$ is a standardized sample (i.e.
-mean 0, variance 1) from the random variable $z_{m}$, thus it equals the sample correlation $\theta_{m}=\sum_{n}x_{n,m}y_{n}=\hat{\rho}_{z_{m},y}$.
-Now we develop the formal demonstration of Proposition 2. We note to the right side a hint to the criteria that was applied to derive every equation.
-$\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}$ $\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$ Definition of $\alpha$. (16)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{E}\left[(z_{m}-\bar{z})^{2}\right]$ Properties of expectations. (17)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{E}[z_{m}-\bar{z}]^{2}+\mathbb{V}[z_{m}-\bar{z}]\right)$ Expectation of a squared variable. (18)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{V}[z_{m}-\bar{z}]$ Means of $z_{m}$ and $\bar{z}$ are 0. (19)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{V}[z_{m}]-\mathbb{V}[\bar{z}]\right)$ Property of sum of variances. (20)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}[\bar{z}]\right)$ Variance of $z_{m}$ is 1. (21)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}\left[\sum_{m^{\prime}}\theta_{m^{\prime}}z_{m^{\prime}}\right]\right)$ Definition of $\bar{z}$. (22)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\mathbb{V}\left[z_{m^{\prime}}\right]\right)$ Properties of variance. (23)
-$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Variance of $z_{m}$ is 1. (24)
-$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Property of sum of limits. (25)
-$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\cdot\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Product-sum property of limits. (26)
-$=\left[\sum_{m}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m},y}\right]$ $\cdot\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given that $\theta_{m}=\hat{\rho}_{z_{m},y}\approx\rho_{z_{m},y}$.
-(27)
-=1=1$\cdot$= 1 ⋅ $\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given Proposition 1. (28)
-$=1-$ $\sum_{m^{\prime}}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m^{\prime}},y}^{2}$ Properties of limits. (29)
-$\lim_{\rho_{z_{p},y}\to 1}$ $\alpha=0$ Given Proposition 1. (30)
-Lemma 1.
-When using dropout, it still holds that $\theta_{m}=\hat{\rho}_{z_{m},y}.$
-Proof.
-$\min_{\Theta}\mathbb{E}_{R\sim\mathrm{Ber(\gamma)}}||Y-(R*X)\Theta||^{2}$ (31)
-$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Gamma\Theta||^{2},$ (32)
-where $\Gamma=(\mathrm{diag(X^{T}X)})^{1/2}$. If we use the uncorrelation of the base models, we obtain:
-$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2}$ (33)
-$=\min_{\Theta}||Y-\tilde{X}\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2},$ (34)
-which is equivalent to ridge regression, with closed form solution:
-$\Theta=$ $\left({\tilde{X}}^{T}{\tilde{X}}+\gamma(1-\gamma)I\right)^{-1}{\tilde{X}}^{T}Y$ (35)
-$=$ $\left([\gamma^{2}+\gamma(1-\gamma)]I\right)^{-1}{\tilde{X}}^{T}Y$ (36)
-$=$ $\left(\gamma I\right)^{-1}{\tilde{X}}^{T}Y$ (37)
-$=$ $\frac{1}{\gamma}{\tilde{X}}^{T}Y$ (38)
-$=$ $X^{T}Y.$ (39)
-Proposition 6 (Avoiding Diversity Collapse).
+Proposition A.1 (Diversity Lower Bound).
 As the correlation of the preferred model $\rho_{p_{m},y}\rightarrow 1$, the diversity $\alpha\rightarrow 1-\gamma$, when using Base Models’ DropOut with probability of retaining $\gamma$.
 Proof.
-We follow a similar procedure as for Proposition 2, by considering $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$.
-We demonstrate that $\mathbb{V}(r\cdot z_{m})=\gamma$, given that $r\sim\mathrm{Bernoulli}(\gamma)$.
-$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})\cdot\mathbb{V}(z_{m})+\mathbb{V}(z_{m})\cdot\mathbb{E}(r_{m})^{2}+\mathbb{V}(r_{m})\cdot\mathbb{E}(z_{m})^{2}$ (40)
-$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})+\mathbb{E}(r_{m})^{2}$ (41)
-$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma(1-\gamma)+\gamma^{2}$ (42)
-$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma.$ (43)
-Then, we evaluate the variance of the ensemble using DropOut $\mathbb{V}(\bar{z})$:
-$\mathbb{V}(\bar{z})=\mathbb{V}{\left(\sum_{m}r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (44)
-$\mathbb{V}(\bar{z})=\sum_{m}\mathbb{V}{\left(r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (45)
-$\mathbb{V}(\bar{z})=\sum_{m}\theta^{2}_{m}\mathbb{V}{\left(r_{m}\cdot z_{m}\right)}$ (46)
-$\mathbb{V}(\bar{z})=\gamma\sum_{m}\theta^{2}_{m}.$ (47)
-Applying Equation 47 into Equation 22, we obtain:
-$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\gamma_{m}\cdot\theta_{m}(z_{m}-\bar{z})^{2}\right]$ (48)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Applying Equation 15 into Equation LABEL:proof2:l7, we obtain:
+[…]
 $=\lim_{\rho_{z_{p},y}\to 1}\sum_{m}\theta_{m}\left(1-\gamma\sum_{m}\theta^{2}_{m}\right)$ (17)
@@
 $\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=1-\gamma.$ (20)
-Appendix B Limitations and Broader Impact.
+Appendix B Limitations, Broader Impact and Future Work
 While our proposed method offers several advantages for post-hoc ensemble selection, it is important to recognize its limitations.
@@
 However, this could also enhance the results presented in this paper.
+Additionally, some bayesian approaches (McEwen et al.,, 2021) could further increase the robustness to the size of the validation data set.
 Our approach is highly versatile and can be seamlessly integrated into a wide variety of ensemble-based learning systems, significantly enhancing their predictive capabilities.
@@
 Instead, its effects will largely depend on how it is applied within different contexts and domains, making its societal implications contingent on the specific use case.
-Appendix C Related Work Addendum
+In the future, we aim to explore in-context learning (Brown et al.,, 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input.
+We discuss broader impact and limitations in Appendix B.
+Appendix C Space complexity of Neural Ensemblers
+[…]
+Appendix D Related Work Addendum
 Ensemble Search via Bayesian Optimization.
-Ensembles of models with different hyperparameters can be built using Bayesian optimization by iteratively swapping a model inside an ensemble with another one that maximizes the expected improvement (Levesque et al., 2016).
-DivBO (Shen et al., 2022) and subsequent work (Poduval et al., 2024) combine the ensemble’s performance and diversity as a measure for expected improvement.
-Besides Bayesian Optimization, an evolutionary search can find robust ensembles of deep learning models (Zaidi et al., 2021).
+Ensembles of models with different hyperparameters can be built using Bayesian optimization by iteratively swapping a model inside an ensemble with another one that maximizes the expected improvement (Levesque et al.,, 2016).
+DivBO (Shen et al.,, 2022) and subsequent work (Poduval et al.,, 2024) combine the ensemble’s performance and diversity as a measure for expected improvement.
+Besides Bayesian Optimization, an evolutionary search can find robust ensembles of deep learning models (Zaidi et al.,, 2021).
 Although these approaches find optimal ensembles, they can overfit the validation data used for fitting if run for many iterations.
-Appendix D Details On Metadatasets
-We provide general information about the datasets in Table 5.
-Table 5: Metadatasets Information
-Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset
-Nasbench (100) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 100
-Nasbench (1K) Vision NAS, Classification (Dong & Yang, 2020) 3 11000 6000 1K
-QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024) 30 160 160 255
-QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024) 30 1088 1088 203
-FTC Language Finetuning, Classification, Section D.1 6 39751 29957 105
-TabRepo Clas. Tabular Classification  (Salinas & Erickson, 2023) 83 1134 126 1530
-TabRepo Reg. Tabular Regression  (Salinas & Erickson, 2023) 17 3054 3397 1530
-Sk-Learn Pipelines. Tabular Classification, Section D.2 69 1514 1514 500
-D.1 Finetuning Text Classifiers (FTC) Metadataset
-Table 6: Search Space.
-Hyperparameter Values
-Model GPT2, Bert-Large, Albert-Large, Bart-Large, T5-Large
-Learning Rate 0.00001, 0.0001, 0.0005, 0.001, 0.005
-LoRA Rank 8, 16, 32, 64, 128
-Search Space
-It comprises three hyperparameters: model, learning rate and LoRA rank (Hu et al., 2022).
-We consider consider five models: 1) GPT2, 124M parameters; (Radford et al., 2019); 2) Bert-Large, 336M ; (Devlin et al., 2018); 3) Bart-Large 400 M, parameters  (Lewis et al., 2019); 4) Albert-Large, 17M parameters  (Lan et al., 2019); and 5) T5-Large, 770 M parameters  (Raffel et al., 2020).
-For the other two hyperparameters we also consider five different values as specified in Table 6.
-Datasets
-The metadataset contains predictions of models finetuned in five metadatasets for text classification: 1) IMDB (Maas et al., 2011); 2) Tweet (Maggie, 2020), 3) News (Zhang et al., 2015), 4) DBpedia (Zhang et al., 2015), 5) SST2 (Socher et al., 2013) and 6) SetFit (Tunstall et al., 2021).
-For every dataset, we create two versions.
-The first one is trained with the complete training data, while the second version is only with a subset of 10 % of the samples.
-All the datasets are for text classification beween 2 to 14 classes, including diverse domains such as movies, reviews, news, tweets, and text entailment data.
-We provide further information about the datasets in Table 7.
-Figure 7: Mean error across datasets.
-Metadataset Creation and Composition
-We created the dataset by finetuning every model to the train split and, subsequently, saving their predictions on the validation and test split.
-The validation split corresponds to 20 % of the available train data.
-As for some datasets (SST-2 and SetFit) the test data is not completely provided by the creators, or it has hidden labels, we obtain it by using 20 % of the remaining test data.
-Specifically The models are fitted up to 5 epochs, using 1 GPU A100 with batch size equal to 2 and no LoRA dropout.
-We vary only the model, learning rate, and LoRA rank, while keeping the default hyperparameters in the Trainer object from the Transformers Library (version 4.41.0).
-Although we evaluate the models in a grid, some runs yielded out-of-memory errors for some configurations.
-In total, the metadataset contains 1134 evaluated configurations, representing around 3800 GPU hours.
-Additionally, we report information about the metadataset in Table 7.
-Table 7: Metadataset Information
-Dataset # Classes # Train Samples # Val. Samples # Test Samples # Confs (100%) # Confs. (10%)
-IMDB (Maas et al., 2011) 2 20,000 5,000 25,000 125 125
-Tweet (Maggie, 2020) 3 27,485 5,497 3,534 100 100
-News (Zhang et al., 2015) 4 96,000 24,000 7,600 99 120
-DBpedia (Zhang et al., 2015) 14 448,000 112,000 70,000 25 65
-SST-2  (Socher et al., 2013) 2 43,103 13,470 10,776 125 125
-SetFit  (Tunstall et al., 2021) 3 393,116 78,541 62,833 25 100
-Hyperparameter Importance
-We explore the importance of two hyperparameters, learning rate, and LoRA rank, by plotting the mean error as a heatmap in Figure 7.
-The error corresponds to the average across different models and datasets.
-The learning rate is an important hyperparameter, while increasing the LoRA rank does not affect the performance significantly in low learning rates.
-This behaviour is interesting, as it showcases that a small rank is enough for a successful finetuning in this context.
-D.2 Scikit learn Pipelines
+Appendix E Baselines Details
+We describe the setup for the baselines evaluation.
+Most of the baselines, except random, use the validation set for training the ensembling model or choosing the base models.
+For all the models that expected a fixed set of $N$ base models, we used $N=50$, following previous work insights (Feurer et al.,, 2015).
+Some metadatasets (Nasbench201-Mini,QuickTune-Mini incurred in a large memory cost that made a few baselines fail in some datasets.
+Since this only happened in less than $10\%$ of the datasets, we imputed the metrics for this specific cases, by using the metric of single-best as the imputation value.
+This mimics a real-world application scenario where the ensembler fall backs to just a single-best approach, which uses few memory for ensembling.
+  • Single best selects the best model according to the validation metric;
+  • Random chooses 50 random models to ensemble;
+  • Top-N ensembles the best $N\in\{5,50\}$ models according to the validation metrics;
+  • Greedy (Caruana et al.,, 2004) creates an ensemble with $N=50$ models by iterative selecting the one that improves the metric as proposed by previous work;
+  • Quick builds the ensemble with 50 models by adding model subsequently only if they strictly improve the metric;
+  • CMAES (Purucker and Beel, 2023b, ) uses an evolutive strategy with a post-processing method for ensembling. We used the implementation in the Phem library (Purucker,, 2024);
+  • Model Average computed the weighted sum of the predictions as in Equation 3. We learn the weight using gradient-descent for optimizing the validation metric;
+  • DivBO (Shen et al.,, 2022): uses Bayesian Optimization for querying the best ensemble, by applying an acquisition function that accounts for both the diversity and performance.
+We implemented the method following the setup described by the authors in their paper;
+  • Ensemble Optimization (Levesque et al.,, 2016): uses Bayesian Optimization for selecting iteratively the model to replace another model inside the ensemble.
+We implemented the method following the default setup described by the authors;
+  • Machine Learning Models as Stackers: We used the default configurations provided by Sckit-learn (Pedregosa et al.,, 2011) for these stackers.
+The input to the models is the concatenation of all the base models’ predictions.
+We concatenated the probabilistic predictions from all the classes in the classification tasks.
+This sometimes produced a large dimensional input space, and a large memory load.
+The model was trained on the validation set.
+Specifically, we apply popular models: SVM, Random Forest, Gradient Boosting, Linear Regression, XGBoost, LightGBM;
+  • Akaike Weighting or Pseudo Model Averaging (Wagenmakers and Farrell,, 2004) computes weights using the relative performance of every model.
+Assuming that the lowest metric from the pool of base models is $\ell_{\text{min}}\in\{\ell_{1},...,\ell_{M}\}$, then the weight is computed using $\Delta_{i}=\ell_{i}-\ell_{\text{min}}$ as:
+    $w_{i}=\frac{\exp^{-\frac{1}{2}\Delta_{i}}}{\sum_{m}\exp^{-\frac{1}{2}\Delta_{m}}}$ (21)
+  • Dynamic Ensemble Search Baselines learn a model that ensembles different models per instance.
+We used baselines from the DESlib library (Cruz et al.,, 2020) such as KNOP (Cavalin et al.,, 2013), KNORAE (Ko et al.,, 2008) and MetaDES (Cruz et al.,, 2015).
+Appendix F Additional Results Related to Research Question #1
+In this Section we report additional results from our experiments:
+  • Average Ranking of baselines for the Negative Log-likelihood (Table 3) and Classification Errors (Table 4).
+  • Average NLL in Scikit-learn Pipelines metadataset (Figure 6).
+Table 3: Average Ranked NLL
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
+Single-Best 17.6667±0.9832 16.6667±2.3094 15.3333±2.7538 7.0167±2.5103 7.5333±2.5221 8.6325±4.6487 9.0294±4.5705
+Random 20.0000±0.0000 9.3333±5.6862 12.3333±2.3094 19.0333±1.3060 19.0167±0.9143 14.0301±5.0166 16.4118±2.9803
+Top5 13.6667±3.2660 11.6667±1.5275 8.3333±1.5275 6.9000±1.9360 8.5667±2.3146 7.0542±4.1083 8.4118±5.1455
+Top50 13.3333±1.7512 5.3333±1.5275 5.0000±1.0000 13.9667±2.0212 13.9500±2.0776 7.6747±3.6729 8.1176±4.4984
+Quick 7.6667±1.9664 7.3333±3.7859 4.3333±3.5119 4.7333±2.0331 6.1333±2.7131 6.9036±3.4273 7.0000±4.1231
+Greedy 4.5000±1.3784 4.3333±3.2146 12.1667±2.7538 3.5167±1.7786 4.7000±2.5278 6.6506±3.6138 7.8824±4.1515
+CMAES 15.3333±5.5737 16.6667±2.3094 15.3333±2.7538 15.8667±4.0809 17.4000±2.1066 11.4578±4.7094 7.3529±2.9356
+Random Forest 8.6667±2.7325 15.6667±3.0551 17.0000±1.7321 15.0000±1.5702 11.1500±4.9674 13.3614±6.1159 9.6471±5.5895
+Gradient Boosting 4.0000±5.4037 17.3333±3.0551 15.8333±3.6171 11.9667±6.8857 12.3000±6.8778 13.2771±5.8400 9.6471±5.4076
+SVM 13.3333±1.6330 11.6667±8.5049 13.0000±8.6603 17.5333±1.5533 14.4500±6.5828 12.4639±5.0827 14.5882±6.9826
+Linear 9.3333±2.1602 13.0000±2.6458 13.0000±3.6056 7.0000±3.1073 6.7833±3.4433 11.8434±6.0897 17.8824±4.4565
+XGBoost 13.1667±5.4559 12.6667±4.1633 14.0000±6.2450 14.1333±2.1413 12.2500±3.0562 16.9819±3.4698 16.1765±3.7953
+CatBoost 3.5000±3.3317 15.0000±2.6458 15.6667±2.3094 11.7000±1.8919 12.9667±3.5548 11.3012±4.8583 10.4118±4.8484
+LightGBM 10.6667±6.1860 17.0000±5.1962 17.0000±5.1962 13.4333±2.8093 15.1333±2.8945 17.6867±3.7900 13.2353±5.2978
+Akaike 13.5000±3.3166 5.0000±1.0000 4.6667±0.5774 12.5333±2.0083 12.6833±1.9761 7.9639±4.0166 7.3824±4.6788
+MA 14.1667±3.8687 8.0000±1.7321 5.0000±3.4641 16.7667±1.4003 16.1000±2.1270 10.3916±4.9226 9.9412±6.5141
+DivBO 8.5000±4.5497 11.0000±6.2450 4.3333±4.0415 4.9500±2.5876 5.2000±2.7687 7.4337±3.5585 7.8824±3.6552
+EO 7.1667±4.5789 6.0000±2.0000 9.0000±2.6458 5.8833±2.2194 5.6167±2.5484 7.0241±3.4144 8.2353±3.4192
+NE-Stack 7.6667±5.9889 1.0000±0.0000 4.0000±5.1962 3.3000±3.6874 2.4000±2.3282 11.7470±7.1326 15.2353±4.6169
+NE-MA 4.1667±3.3116 5.3333±3.5119 4.6667±2.3094 4.7667±2.1445 5.6667±2.1389 6.1205±3.8553 5.5294±2.7184
+Table 4: Average Ranked Error
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
+Single-Best 14.6667±3.4593 18.3333±0.2887 17.1667±1.5275 11.0167±4.8128 9.7333±4.0231 10.7048±5.2452 11.2840±4.3236
+Random 19.3333±1.6330 14.6667±1.1547 14.8333±0.7638 19.8667±0.3198 19.6667±0.6065 13.8193±6.5698 15.0494±5.8130
+Top5 6.5833±4.4768 13.3333±2.0817 10.6667±3.5119 4.8833±3.3725 4.8667±3.3859 10.5181±4.6751 8.5741±5.0013
+Top50 16.5000±3.3317 4.3333±1.5275 2.3333±1.5275 9.6333±4.9444 10.9000±3.2094 9.6988±5.1234 8.6605±4.7189
+Quick 6.8333±4.8442 8.6667±1.1547 7.3333±2.3094 6.1000±4.0480 3.6333±2.4066 10.0301±4.0412 7.9136±4.9319
+Greedy 5.5833±4.3865 13.0000±3.4641 11.3333±1.5275 10.1167±4.3543 7.5000±3.7093 9.7108±4.1069 8.1975±4.5830
+CMAES 5.0833±2.3752 6.0000±3.6056 7.3333±2.0817 9.6333±3.9105 6.6167±2.6152 10.5060±4.8341 11.6173±6.0220
+Random Forest 7.1667±4.6224 9.0000±5.1962 11.1667±5.5752 13.8000±4.5837 14.3667±2.7697 10.5542±4.6302 10.6667±5.9119
+Gradient Boosting 12.5000±4.4159 19.0000±0.8660 17.6667±2.2546 9.6667±5.1551 14.2333±4.1351 11.0181±5.9671 11.0185±6.6635
+SVM 3.8333±3.3714 9.0000±1.7321 10.5000±5.8949 12.0000±6.0955 15.6500±3.5261 8.3855±4.7788 13.0185±7.1144
+Linear 8.8333±3.1252 13.6667±5.0332 14.1667±3.6856 6.2500±3.5834 7.9333±3.2872 9.3916±6.3745 10.9259±6.7519
+XGBoost 11.4167±5.5355 14.6667±2.3094 13.0000±1.7321 17.0000±3.2350 16.1333±3.0932 12.5120±5.8304 11.6111±4.4651
+CatBoost 14.6667±3.3267 17.0000±0.0000 17.5000±1.5000 12.0500±3.8155 14.5167±3.2310 11.9699±5.5179 11.4506±4.2834
+LightGBM 10.5833±5.0241 17.0000±5.1962 17.0000±5.1962 15.7833±3.4433 17.5167±2.4193 11.7470±6.1947 11.4815±4.6094
+Akaike 13.3333±6.6608 2.6667±2.0817 3.3333±0.5774 9.1500±4.8551 10.1500±3.5016 9.9639±5.0677 11.7654±4.2925
+MA 16.0000±3.4641 5.3333±1.1547 2.6667±2.0817 11.6833±4.3519 12.7833±3.7132 10.1024±5.3487 10.4815±4.7994
+DivBO 11.2500±4.8760 12.3333±1.1547 15.6667±2.7538 11.2167±3.8117 7.9000±3.8336 9.8193±4.5830 10.1667±4.6657
+EO 12.6667±5.7504 5.3333±3.7859 4.6667±2.5166 10.5167±3.4851 9.0667±3.0618 9.9036±5.3905 8.8889±4.0226
+NE-Stack 5.5000±3.3317 4.0000±1.7321 8.3333±0.5774 3.7500±2.7189 2.7833±2.2995 10.8313±6.6403 9.3148±7.0986
+NE-MA 7.6667±2.2509 2.6667±1.5275 3.3333±2.5166 5.8833±4.3225 4.0500±2.4046 8.8133±4.9137 7.9136±4.8456
+Figure 6: Average Normalized NLL in Scikit-learn pipelines (Validation).
+F.1  Unnormalized Results
+We report the results for the research question 1 in Tables 5 and 6, omitting the normalization.
+This helps us to understand how much the normalization is changing the results.
+However, as the different datasets have metrics with different scales, it is important to normalize the results to get a better picture of the relative performances.
+This is especially problematic in the regression tasks, as it depends on the target scale, therefore we omit it.
+Even without the normalization, our proposed approach achieves the best results across different datasets.
+The ranking results remain the same independently of the normalization.
+Table 5: Average Unnormalized NLL.
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class
+Single-Best 0.3412±0.3043 1.9175±1.2203 1.5696±0.8119 0.4499±0.5107 0.6339±0.5276 0.3647±0.3300
+Random 0.4521±0.2982 1.2202±1.0069 1.3005±0.9988 2.1473±0.6480 2.7139±1.0686 0.4344±0.3341
+Top5 0.2975±0.2799 1.2940±0.9934 1.1911±0.9833 0.4234±0.4473 0.6702±0.5330 0.3576±0.3306
+Top50 0.2819±0.2631 1.1515±0.9713 1.1496±0.9797 0.7813±0.5492 1.2587±0.8605 0.3561±0.3331
+Quick 0.2443±0.2128 1.1679±0.9614 1.1455±0.9608 0.3952±0.4441 0.6030±0.4869 0.3562±0.3326
+Greedy 0.2366±0.2107 1.1498±0.9652 1.2953±0.9657 0.3895±0.4453 0.5817±0.4870 0.3566±0.3317
+CMAES 0.3552±0.2130 1.9175±1.2203 1.5696±0.8119 1.1665±0.7614 1.8487±0.6070 0.3862±0.3325
+Random Forest 0.2434±0.2033 1.7354±1.4580 1.6589±1.4265 0.8846±0.6550 0.9870±0.8090 0.4452±0.3734
+Gradient Boosting 0.2262±0.1915 2.2979±0.5750 1.8206±0.6327 1.2480±1.6045 1.7043±1.9723 0.4708±0.4264
+SVM 0.2626±0.2241 1.6698±1.4024 1.6590±1.4274 1.3608±0.3953 1.8931±1.2176 0.3951±0.3356
+Linear 0.2496±0.2130 1.3021±0.9815 1.3626±0.9836 0.4701±0.5151 0.6582±0.6088 0.4095±0.3914
+XGBoost 0.2678±0.2250 1.5484±1.2487 1.5788±1.2796 0.8328±0.4807 0.9423±0.5402 0.5503±0.4738
+CatBoost 0.2329±0.2074 1.6813±1.3589 1.6531±1.3327 0.6392±0.6006 1.0816±0.7100 0.3954±0.3839
+LightGBM 0.2442±0.2000 6.7760±5.5841 6.7783±5.5802 0.7957±0.7383 2.5903±3.4557 0.6301±0.5723
+Akaike 0.3012±0.3010 1.1503±0.9699 1.1499±0.9822 0.7369±0.5221 1.0655±0.6000 0.3561±0.3334
+MA 0.2804±0.2074 1.1612±0.9664 1.1530±0.9720 1.0974±0.5102 1.5240±0.8090 0.3788±0.3342
+DivBO 0.2643±0.2526 1.2355±0.8746 1.1659±0.9729 0.4065±0.4489 0.5906±0.4825 0.3588±0.3339
+EO 0.2599±0.2481 1.1532±0.9801 1.2024±0.9656 0.4144±0.4542 0.5937±0.4804 0.3572±0.3329
+NE-Stack 0.2375±0.2031 1.0706±0.9554 1.1543±1.0696 0.3747±0.4494 0.4923±0.4658 0.4587±0.4479
+NE-MA 0.2399±0.2176 1.1486±0.9892 1.1596±0.9991 0.3998±0.4474 0.5832±0.4900 0.3536±0.3311
+Table 6: Average Unnormalized Error.
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
+Single-Best 0.0868±0.0853 0.4320±0.3331 0.4509±0.3019 0.1285±0.1499 0.1667±0.1477 0.1604±0.1545 0.1169±0.1164
+Random 0.1012±0.0910 0.3092±0.2444 0.3502±0.2417 0.5677±0.2487 0.5295±0.2306 0.1840±0.1632 0.1422±0.1356
+Top5 0.0831±0.0825 0.3005±0.2374 0.3007±0.2251 0.1079±0.1375 0.1502±0.1359 0.1594±0.1527 0.1125±0.1142
+Top50 0.0881±0.0829 0.2778±0.2266 0.2788±0.2262 0.1334±0.1495 0.1816±0.1555 0.1584±0.1524 0.1139±0.1160
+Quick 0.0831±0.0824 0.2840±0.2304 0.2842±0.2289 0.1121±0.1409 0.1504±0.1425 0.1583±0.1517 0.1145±0.1192
+Greedy 0.0824±0.0811 0.3253±0.2827 0.2958±0.2368 0.1251±0.1487 0.1596±0.1450 0.1562±0.1523 0.1140±0.1186
+CMAES 0.0822±0.0810 0.2788±0.2247 0.2869±0.2358 0.1234±0.1478 0.1579±0.1456 0.1589±0.1541 0.1269±0.1311
+Random Forest 0.0833±0.0829 0.2958±0.2444 0.3706±0.3698 0.1666±0.1955 0.2122±0.1874 0.1591±0.1528 0.1198±0.1210
+Gradient Boosting 0.0838±0.0809 0.4682±0.2852 0.4920±0.2559 0.1425±0.1773 0.2092±0.1855 0.1620±0.1592 0.1240±0.1293
+SVM 0.0819±0.0806 0.2917±0.2431 0.3356±0.2444 0.1627±0.2031 0.2341±0.2032 0.1553±0.1550 0.1407±0.1358
+Linear 0.0830±0.0816 0.3098±0.2200 0.3469±0.2097 0.1191±0.1503 0.1670±0.1547 0.1577±0.1568 0.1195±0.1203
+XGBoost 0.0848±0.0840 0.3546±0.2824 0.3554±0.2836 0.1903±0.1452 0.2144±0.1598 0.1654±0.1574 0.1193±0.1199
+CatBoost 0.0874±0.0863 0.4022±0.3034 0.4217±0.3411 0.1450±0.1754 0.2090±0.1822 0.1629±0.1606 0.1199±0.1224
+LightGBM 0.0843±0.0833 0.5941±0.4476 0.5907±0.4475 0.1666±0.1604 0.3442±0.3146 0.1633±0.1586 0.1196±0.1222
+Akaike 0.0857±0.0802 0.2767±0.2252 0.2797±0.2273 0.1311±0.1487 0.1769±0.1538 0.1589±0.1527 0.1228±0.1305
+MA 0.0874±0.0803 0.2795±0.2271 0.2765±0.2248 0.1435±0.1618 0.1971±0.1651 0.1609±0.1580 0.1186±0.1262
+DivBO 0.0846±0.0832 0.3024±0.2489 0.3817±0.1893 0.1295±0.1431 0.1616±0.1458 0.1573±0.1540 0.1170±0.1200
+EO 0.0851±0.0829 0.2786±0.2283 0.2813±0.2274 0.1330±0.1561 0.1659±0.1471 0.1589±0.1577 0.1177±0.1237
+NE-Stack 0.0824±0.0817 0.2781±0.2275 0.2879±0.2339 0.1033±0.1348 0.1424±0.1314 0.1607±0.1511 0.1152±0.1159
+NE-MA 0.0828±0.0818 0.2772±0.2264 0.2773±0.2262 0.1093±0.1344 0.1490±0.1383 0.1551±0.1515 0.1149±0.1211
+Figure 7: Critical Difference (CD) diagrams aggregating datasets across metadatasets
+Figure 8: Critical Difference (CD) diagrams aggregating datasets across metadatasets
+Appendix G Critical Difference Diagrams
+We want to gain deeper insights into the difference between our proposed method and the baselines.
+Critical Difference Diagrams.
+To this end, we present critical difference (CD) diagrams to visualize and statistically analyze the performance of different methods across multiple datasets.
+The CD diagrams are generated using the autorank library [^3] https://github.com/sherbold/autorank, which automates the statistical comparison by employing non-parametric tests like Friedman test followed by the Nemenyi post-hoc test.
+These diagrams show the average ranks of the methods along the horizontal axis, where methods are positioned based on their performance across datasets.
+The critical difference value, which depends on the number of datasets, is represented by a horizontal bar above the ranks.
+Methods not connected by this bar exhibit statistically significant differences in performance.
+Evaluation.
+We computed CD diagrams across all datasets in all metadatasets (Figures 7(a)-8(a)) and an aggregated diagram across all datasets in all metadatasets (Figure 8(b)).
+We observe that although the performance difference is not substantial compared to other top-performing post-hoc ensembles like Greedy and Quick, our Neural Ensembler (NE-MA) consistently achieves the best performance in the aggregated results (Figure 8(b)).
+We also highlight that the NE versions were the top-performing approaches across all metadatasets except FTC, even though we did not modify the method’s hyperparameters.
+Significance on Challenging Datasets.
+Given the high performance between Greedy and NE-MA, we wanted to understand when the second one would obtain strong significant results.
+We found that that NE-MA is particularly well performing in challenging datasets with a large number of classes.
+Given Table 11, we can see that four meta-datasets have a high ($>10$) number of classes, thus they have datasets with a lot of classes.
+We selected these metadatasets (NB(100), NB(1000), QT-Micro, QT-Mini), and plotted the significance compared to Greedy and Random Search.
+The results reported in Figure 9 demonstrate that our approach is significantly better than Greedy in these metadatasets.
+Figure 9: CD Diagram in datasets with a lot of classes.
+Appendix H Empirical Analysis of Computational Cost
+To better understand the trade-off between performance and computational cost, we analyze the average normalized Negative Log-Likelihood (NLL) and the runtime of various ensembling techniques.
+Our focus is on post-hoc ensembling methods, assuming that all base models are pre-trained.
+This allows us to isolate and compare the efficiency of the ensembling processes themselves, in contrast to methods like DivBO, which sequentially train models during the search, leading to higher computational demands.
+We measure only the post-hoc ensembling time, i.e.
+training the ensembler or selecting the group of base models for the final ensemble.
+As shown in Figure 10, the Neural Ensemblers (NE-MA and NE-Stack) achieve the best average performance while maintaining a competitive runtime.
+Notably, our method has a shorter runtime than the Greedy ensembling method and surpasses it in terms of performance.
+Additionally, the Neural Ensemblers are faster than traditional machine learning models such as Gradient Boosting, Support Vector Machines (SVM), Random Forests, and optimization algorithms like CMAES.
+While simpler methods like Top5 and Top50 exhibit faster runtimes, they do so at the expense of reduced accuracy.
+Figure 10: Trade-off between performance and computational cost for different ensembling methods.
+The Neural Ensemblers achieve the best performance with competitive runtime, outperforming other methods in both accuracy and efficiency.
+Appendix I Proof-of-Concept with Overparameterized Base Models
+A key question is whether dynamic ensemblers like our Neural Ensembler (NE) offer benefits when base models are overparameterized and potentially overfit the data.
+Specifically, does the advantage of the NE persist in scenarios where the base models have high capacity?
+To explore this, we extend our Proof-of-Concept experiment by ensembling 10th-degree polynomials instead of 2nd-degree ones, thereby increasing the complexity of the base models and introducing the risk of overfitting.
+We follow the same protocol as in Section 2.1, comparing the performance of the NE with the fixed-weight ensemble method.
+As illustrated in Figure 11, even with overparameterized base models that tend to overfit the training data, the NE (specifically the weighted Model-Averaging version) achieves the best performance on unseen data.
+This improvement occurs because the NE dynamically adjusts the ensemble weights based on the input.
+Thus, dynamic ensembling is advantageous not only when base models underfit but also when they overfit the data.
+In contrast, the fixed-weight approach lacks this adaptability and cannot compensate for the overfitting behavior of the base models.
+Figure 11: Proof-of-Concept experiment with overparameterized base models (10th-degree polynomials).
+The Neural Ensembler outperforms fixed-weight ensemble method by mitigating overfitting through dynamic, input-dependent weighting.
+Appendix J Ablation Study on Validation Data Size
+To assess the Neural Ensembler’s (NE) dependence on validation data size and its sample efficiency, we conducted an ablation study by varying the proportion of validation data used for training.
+In this study, we evaluate the NE in stacking mode across all metadatasets, using different percentages of the available validation data: $1\%,5\%,10\%,25\%,50\%,100\%$.
+For each configuration, we compute the Negative Log Likelihood (NLL) on the test set and normalize these values by the performance achieved when using $100\%$ of the validation data.
+This normalization allows us to compare performance changes across different tasks, accounting for differences in metric scales.
+A normalized NLL value below 1 indicates performance better than the baseline with full validation data, while a value above 1 indicates a performance drop.
+The results are presented in Figure 12.
+We observe that reducing the amount of validation data used for training the NE leads to a relative degradation of the performance.
+However, the performance drops are relatively modest in three metadatasets.
+For these experiments we used the same dropout rate 0.75.
+We could improve the robustness in lower percentages of validation data by using a higher dropout rate.
+The DropOut mechanism prevents overfitting by randomly omitting base models during training, while parameter sharing reduces the number of parameters and promotes learning common representations.
+Figure 12: Ablation study on the percentage of validation data used for training the Neural Ensembler.
+The normalized NLL is plotted against the percentage of validation data, with values below 1 indicating performance better than or equal to using the full validation set.
+Appendix K Effect of Merging Training and Validation Data
+In our experimental setup (Section 4), we train the base models using the training split and the ensemblers using the validation split, then evaluate on the test split.
+An important question is whether merging the training and validation data could improve the performance of both the base models and the ensemblers.
+Specifically, we explore:
+  1. (a)
+    Can baseline methods that do not require a validation split, such as Random, achieve better performance if the base models are trained on the merged dataset (training + validation)?
+  2. (b)
+    Would training both the base models and the ensemblers on the merged dataset be beneficial, given that more data might enhance their learning?
+To investigate these questions, we conducted experiments on two metadatasets: Scikit-learn Pipelines and FTC.
+We trained the base models on the merged dataset and also trained the ensemblers on this same data.
+Five representative baselines were compared: NE-MA, NE-Stack, Greedy, Random and Single-best.
+Figure 13(b) presents the test set performance on the Scikit-learn Pipelines metadataset.
+Training on the merged dataset did not improve performance compared to training on the original splits.
+In fact, the results are similar to the Random ensembling method, indicating no significant gain.
+This suggests that training both the base models and ensemblers on the same (larger) dataset may lead to overfitting, hindering generalization to unseen data.
+Further evidence of overfitting is shown in Figure 13(a), where we report the Negative Log Likelihood (NLL) on the merged dataset (effectively the training data).
+The NLL values are significantly lower for models trained on the merged dataset, confirming that they fit the training data well but do not generalize to the test set.
+Similar observations were made on the FTC metadataset, as presented in Table 14(a) and Figure 14(b).
+Although the Random method trained on the merged dataset shows a slight improvement, the overall performance gains are minimal.
+From these results, we conclude that:
+  1. (a)
+    Training base models on the merged dataset does not significantly enhance the performance of baseline methods that do not require a validation split.
+  2. (b)
+    Using the merged dataset to train both the base models and the ensemblers is not beneficial and may lead to overfitting, reducing generalization to the test set.
+(a) Performance on merged dataset (training data) for Scikit-learn Pipelines.
+(b) Test set performance on Scikit-learn Pipelines with merged training.
+Figure 13: Impact of training on merged dataset for Scikit-learn Pipelines.
+Training on the merged dataset leads to overfitting, as evidenced by low NLL on training data but no improvement on test data.
+Algorithm FTC (Avg. Normalized NLL) FTC (Avg. Rank)
+Single-Best $1.0000_{\pm 0.0000}$ $11.4167_{\pm 1.5626}$
+Single-Best (Merged) $1.2816_{\pm 0.6710}$ $11.2500_{\pm 4.0466}$
+Random $1.5450_{\pm 0.5289}$ $14.0000_{\pm 0.8944}$
+Random (Merged) $1.5365_{\pm 0.7142}$ $13.0833_{\pm 2.2454}$
+Top5 $0.8406_{\pm 0.0723}$ $8.6667_{\pm 2.9439}$
+Top50 $0.8250_{\pm 0.1139}$ $8.3333_{\pm 1.9664}$
+Quick $0.7273_{\pm 0.0765}$ $4.6667_{\pm 1.7512}$
+Greedy $\mathbf{0.6943}_{\pm 0.0732}$ $\mathbf{2.8333}_{\pm 1.6021}$
+Greedy (Merged) $0.7537_{\pm 0.2652}$ $6.5833_{\pm 4.3637}$
+CMAES $1.2356_{\pm 0.5295}$ $10.3333_{\pm 3.4448}$
+MA $0.9067_{\pm 0.1809}$ $8.8333_{\pm 2.4014}$
+NE-Stack $0.7562_{\pm 0.1836}$ $5.3333_{\pm 4.6762}$
+NE-Stacking (Merged) $0.7428_{\pm 0.2208}$ $5.5833_{\pm 3.3529}$
+NE-MA $\underline{\mathit{0.6952}}_{\pm 0.0730}$ $\mathbf{2.8333}_{\pm 2.2286}$
+NE-MA (Merged) $0.7461_{\pm 0.2084}$ $6.2500_{\pm 2.6410}$
+Figure 14: Merged-split baselines on the FTC metadataset: (a) Test set performance with merged training, (b) Critical Difference diagram.
+Appendix L Neural Ensemblers Operating on the Original Input Space
+In Section 3.1, we discussed that the Neural Ensembler in model-evaraging mode (NE-MA) computes weights $\theta_{m}(z;\beta)$ that rely solely on the base model predictions $z$ for each instance.
+Specifically, the weights are defined as:
+$\theta_{m}(z;\beta)=\frac{\exp f_{m}(z;\beta)}{\sum_{m^{\prime}}f_{m^{\prime}}(z;\beta)}$ (22)
+where $z$ represents the base model predictions for an instance $x\in\mathcal{X}$, and $\mathcal{X}$ denotes the original input space.
+As our experiments encompass different data modalities, $x$ can be a vector of tabular descriptors, an image, or text.
+An alternative formulation involves computing the ensemble weights directly from the original input instances instead of using the base model predictions.
+This approach modifies the weight computation to:
+$\hat{y}=\sum_{m}\theta_{m}(x;\beta)\cdot z_{m}(x)=\sum_{m}\frac{\exp f_{m}(x;\beta)}{\sum_{m^{\prime}}\exp f_{m^{\prime}}(x;\beta)}\cdot z_{m}(x)$ (23)
+where $f_{m}$ now operates on the original input space $\mathcal{X}$ to produce unnormalized weights.
+However, adopting this formulation introduces challenges in selecting an appropriate function $f_{m}$ for different data modalities.
+For example, if the instances are images, $f_{m}$ must be a network capable of processing images, such as a convolutional neural network.
+This requirement prevents us from using the same architecture across all modalities, limiting the generalizability of the approach.
+To evaluate this idea, we tested this alternative neural ensembler on the Scikit-learn Pipelines metadataset, which consists of tabular data.
+We implemented $f_{m}$ as a four-layer MLP.
+Our results, represented by the yellow bar in Figure 13(b), indicate that this strategy does not outperform the original approach proposed in Section 3, which uses the base model predictions as input.
+We hypothesize that computing ensemble weights directly from the original input space may be more susceptible to overfitting, especially when dealing with datasets that have noisy or high-dimensional features.
+Additionally, this strategy may require tuning the hyperparameters of the network $f_{m}$ for each dataset to achieve optimal performance, reducing its effectiveness and generalizability across diverse datasets.
+Appendix M Do Neural Ensemblers need a strong group of base models, i.e. found using Bayesian Optimization?
+Experimental Protocol.
+Practitioners use some methods such as greedy ensembling as post-hoc ensemblers, i.e., they consider a set of models selected by a search algorithm such as Bayesian Optimization as base learners.
+DivBO enhances the Bayesian Optimization by accounting for the diversity in the ensemble in the acquisition function.
+We run experiments to understand whether the Neural Ensemblers’ performance depends on a strong subset of $50$ base models selected by DivBO, and whether it can help other methods.
+We conduct additional experiments by randomly selecting $50$ models to understand the impact and significance of merely using a smaller set of base models.
+We normalize the base of the metric on the single-best base model from the complete set contained in the respective dataset.
+Results.
+We report in Table 7 the results with the two selection methods (random and DivBO) using a subset of common baselines, where we normalize using the metric of the single-best from the whole set of models.
+We can compare directly with the results in Table 1.
+We observe that reducing the number of base models with DivBO negatively affects the performance of the Neural Ensemblers.
+Surprisingly, randomly selecting the subset of base models improves the results in two metadatasets (TR-Class and NB-1000).
+We hypothesize that decreasing the number of base models is beneficial for these metadatasets.
+With over $1000$ base models available, the likelihood of identifying a preferred model and overfitting the validation data increases in these metadatasets.
+Naturally, decreasing the number of base models can also be detrimental for the Neural Ensemblers, as this happens for some metadatasets such as TR-Reg and QT-Micro.
+In contrast to the Neural Ensemblers, selecting a subset of strong models with DivBO improves the performance for some baselines such as Model Averaging (MA) or TopK ($K=25$).
+In other words, it works as a preprocessing method for these ensembling approaches.
+Overall, the results in Table 7 demonstrate that Neural Ensemblers do not need a strong group of base models to achieve competitive results.
+Table 7: Average NLL for Subset of Base Models selected Randomly or using DivBO
+FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
+Single 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
+Single + DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
+Random + DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
+Top25 + DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
+Quick + DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
+Greedy + DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
+CMAES + DivBO 0.8915±0.1759 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0166±0.3319 1.0265±0.1521
+Random Forest + DivBO 0.7932±0.1194 0.9338±0.3436 1.1609±0.2787 2.5998±2.6475 1.7330±0.9898 1.3308±0.6640 1.0559±0.1240
+Gradient Boosting + DivBO 0.7908±0.1848 1.4011±0.5359 1.0000±0.0000 3.1558±1.8843 3.0103±1.3714 1.3173±1.1519 1.1285±0.4153
+Linear + DivBO 0.7433±0.0870 0.6471±0.2272 1.0199±0.0345 1.5843±1.6159 1.2392±0.5343 1.0786±1.0100 1.0326±0.1265
+SVM + DivBO 0.8312±0.0943 0.7406±0.2612 1.0730±0.1264 5.5177±3.3684 5.0079±3.4926 1.4542±1.4518 2.7702±2.9398
+XGB + DivBO 0.7884±0.1286 0.7519±0.2361 1.0000±0.0000 3.1705±2.8221 2.0159±1.8026 1.7698±1.6369 1.3226±0.6368
+CatBoost + DivBO 0.6941±0.0953 0.8110±0.2405 1.0000±0.0000 2.3936±2.4785 2.4792±2.3078 1.3149±1.3954 1.0730±0.0815
+LightGBM + DivBO 0.7706±0.1919 3.7392±2.7136 1.0000±0.0000 2.1979±2.2093 2.2706±2.2088 1.6938±1.1246 1.6427±2.0133
+Akaike + DivBO 0.8757±0.0944 0.8159±0.2196 1.0000±0.0000 1.2270±1.0049 1.0452±0.2362 1.0202±0.3379 1.0564±0.1884
+MA + DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
+Single + Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
+Top25 + Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
+Quick + Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
+Greedy + Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
+CMAES + Random 1.2164±0.3851 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.1579±0.9913 1.0004±0.0851
+RF + Random 0.7505±0.0915 0.8325±0.2255 0.8654±0.3179 3.8684±2.8060 3.0156±1.9975 1.2034±0.4063 1.0079±0.0795
+GBT + Random 0.7235±0.1605 1.8377±1.4510 0.9533±0.0809 2.3407±1.3498 3.1659±2.0795 1.3344±1.1611 1.0500±0.1940
+Linear + Random 0.7541±0.0897 0.7400±0.2827 0.7732±0.2331 2.0502±2.2932 1.9028±1.3239 1.0423±1.0174 1.0430±0.1224
+SVM + Random 0.8010±0.0903 0.7767±0.3006 0.9268±0.5498 5.4773±3.3607 5.5665±3.3068 1.3964±1.4267 2.8271±3.0203
+XGB + Random 0.8288±0.1463 0.7369±0.2384 0.8875±0.5032 3.7747±3.1635 2.6163±2.1103 1.7214±1.5419 1.1962±0.3298
+CatBoost + Random 0.6911±0.1007 0.8294±0.2366 0.9472±0.4823 2.6695±2.6137 2.7315±2.3219 1.2006±1.2116 1.0559±0.2144
+LightGBM + Random 0.7960±0.1977 3.6975±2.6631 5.2689±4.6347 2.8698±2.6137 3.6272±3.2791 1.7133±1.0934 1.6848±2.1630
+Akaike + Random 0.8045±0.0987 0.6538±0.1831 0.6905±0.2981 2.2022±2.4118 1.5073±0.6848 1.1180±1.0381 0.9970±0.0886
+MA + Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
+NE(S) + Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
+NE(MA) + Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
+NE(S) + DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
+NE(MA) + DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
+Figure 15: Mean weights assigned to the base models decrease with DropOut rate. Every data point is a model. The errors and weights are the mean values across many datasets for every model.
+Appendix N Visualizing the effect of the regularization on the model weights
+In Figure 15, we study the effect of the DropOut rate on the ensembler weights, in Model-Averaging model using the QT-Micro meta-dataset.
+When there is no DropOut some weights are close to one, i.e.
+they are preferred models.
+As we increase the value, many models with high weights decrease.
+If the rate is very high (e.g.
+$0.9$), we will have many models contributing to the ensemble, with weights different from zero.
+Appendix O Details On Meta-Datasets
+In Table 11 we provide further information about the meta-datasets, such as the number of datasets, the average samples in the test and validation splits and the average number of classes.
+O.1  Scikit learn Pipelines
 Search Space
@@
 To effectively study ensembling methods and benchmark different strategies, we require a diverse set of pipelines.
-Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al., 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers.
-The pipelines are structured in three stages—preprocessor, feature selector, and classifier—which allows us to systematically explore numerous configurations.
+Therefore, we construct a comprehensive search space inspired by the TPOT library (Olson et al.,, 2016), encompassing a wide range of preprocessors, feature selectors, and classifiers.
+The pipelines are structured in three stages: preprocessor, feature selector, and classifier.
+This allows us to explore numerous configurations systematically.
 This extensive and diverse search space enables us to examine the impact of ensembling on a variety of models and serves as a robust benchmark for evaluating different ensembling techniques.
 Detailed descriptions of the components and their hyperparameters are provided in Tables 8, 9, and 10.
+Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al.,, 2019) as the foundation for our meta-dataset.
+OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains.
+This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
+Meta-Dataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space.
+Each pipeline execution was constrained to a maximum runtime of $15$ minutes.
+During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints.
+For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features.
+We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset.
+This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
 Table 8: Classifiers and their hyperparameters used in the TPOT search space.
@@
    max_features (float, [0.05, 1.0], default=0.525)
-Datasets We utilized the OpenML Curated Classification benchmark suite 2018 (OpenML-CC18) (Bischl et al., 2019) as the foundation for our meta-dataset.
-OpenML-CC18 comprises $72$ diverse classification datasets carefully selected to represent a wide spectrum of real-world problems, varying in size, dimensionality, number of classes, and domains.
-This selection ensures comprehensive coverage across various types of classification tasks, providing a robust platform for evaluating the performance and generalizability of different ensembling approaches.
-Metadataset Creation To construct our meta-dataset, we randomly selected $500$ pipeline configurations for each dataset from our comprehensive search space.
-Each pipeline execution was constrained to a maximum runtime of $15$ minutes.
-During this process, we had to exclude three datasets (connect-4, Devnagari-Script, Internet-Advertisements) due to excessive computational demands that exceeded our runtime constraints.
-For data preprocessing, we standardized the datasets by removing missing values and encoding categorical features.
-We intentionally left other preprocessing tasks to be handled autonomously by the pipelines themselves, allowing them to adapt to the specific characteristics of each dataset.
-This approach ensures that the pipelines can perform necessary transformations such as scaling, normalization, or feature engineering based on their internal configurations, which aligns with our objective of evaluating automated machine learning pipelines in a realistic setting.
-Appendix E Additional Results
-In this Section we report additional results from our experiments:
-  • Average Ranking of baselines for the Negative Log-likelihood (Table 11) and Classification Errors 12.
-  • Average NLL in Scikit-learn pipelines metadataset 8.
-  • Average NLL using a subset of base models selected via DivBO 13, random 14 or both 15.
-Table 11: Average Ranked NLL
-FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Reg
-Single-Best 14.0000±0.8944 14.0000±1.0000 12.8333±1.0408 6.8833±2.3219 7.3667±2.5049 7.5361±3.8531 7.7647±3.9096
-Random 16.0000±0.0000 7.6667±4.0415 9.3333±0.5774 15.2500±0.8068 15.2333±0.7739 12.0000±3.9962 13.6176±1.8669
-Top5 11.3333±2.5033 9.6667±1.1547 7.0000±1.0000 6.9000±1.9360 8.3000±1.9678 6.2470±3.3951 7.2353±4.2907
-Top50 10.8333±1.7224 4.6667±1.1547 4.3333±0.5774 11.4000±1.2959 11.5333±1.5025 6.8554±3.4006 6.8824±3.7397
-Quick 6.1667±1.6021 6.3333±3.0551 3.6667±2.5166 4.7333±2.0331 6.0333±2.5661 6.2048±2.8531 5.9412±3.2107
-Greedy 3.5000±0.8367 4.0000±2.6458 12.8333±1.0408 3.5167±1.7786 4.5667±2.4238 5.9036±3.0268 6.5882±3.2607
-CMAES 12.1667±4.4460 14.0000±1.0000 12.8333±1.0408 12.9667±2.6811 13.9667±1.3322 9.9277±3.8880 6.2353±2.4630
-Random Forest 7.0000±2.7568 13.0000±2.6458 14.0000±2.6458 11.9833±1.2964 9.3000±3.8699 11.4217±5.1493 8.2353±4.8157
-Gradient Boosting 3.1667±4.8339 14.6667±1.5275 13.5000±2.1794 9.7000±5.2203 10.1167±5.1890 11.4217±4.9463 8.3529±4.7558
-SVM 10.6667±1.3663 10.0000±7.0000 11.3333±7.2342 13.9167±1.2532 11.6833±5.0198 10.8434±4.2641 11.7941±5.5707
-Linear 7.6667±2.0656 11.0000±1.0000 10.0000±2.0000 6.8000±2.8816 6.5833±3.3479 10.3855±5.2587 14.3529±3.4989
-MA 11.1667±2.8577 6.6667±1.1547 4.3333±2.3094 13.2167±1.0059 12.8167±1.6000 9.0241±4.0567 8.2941±5.1813
-DivBO 7.1667±3.5449 9.0000±4.5826 5.3333±5.7735 4.9167±2.5123 5.1000±2.7114 6.6145±2.9377 6.7647±2.9054
-EO 5.8333±3.7103 5.3333±1.5275 7.3333±1.5275 5.8833±2.2194 5.5000±2.5052 6.2169±2.6689 7.0000±2.5739
-NE-Stack (Ours) 6.0000±4.7329 1.0000±0.0000 3.3333±4.0415 3.2000±3.4381 2.3333±2.0899 10.0241±5.8351 12.2941±3.5314
-NE-MA (Ours) 3.3333±2.5033 5.0000±3.0000 4.0000±1.7321 4.7333±2.0331 5.5667±1.9945 5.3735±3.2674 4.6471±2.3702
-Table 12: Average Ranked Error
-FTC NB-Micro NB-Mini QT-Micro QT-Mini TR-Class TR-Class (AUC)
-Single-Best 12.1667±2.5626 15.0000±0.8660 13.8333±0.7638 9.8167±3.7908 9.1167±3.5591 8.9157±4.1480 9.3086±4.1030
-Random 15.8333±0.4082 13.0000±0.0000 12.0000±1.0000 15.9667±0.1826 15.9667±0.1826 11.3253±5.0723 12.2716±4.3747
-Top5 6.0833±4.2710 11.6667±2.0817 8.3333±2.0817 4.4000±2.8780 4.6167±2.9204 8.7048±3.6325 7.2099±4.0519
-Top50 13.5000±2.2583 3.3333±1.5275 2.0000±1.0000 8.3000±4.1618 9.7333±2.6546 7.9880±4.1548 7.2222±3.6929
-Quick 6.3333±4.1433 7.6667±1.1547 6.3333±2.3094 5.5833±3.5186 3.5500±2.3575 8.3253±3.1771 6.8148±3.8033
-Greedy 5.0833±3.8264 11.3333±3.0551 13.8333±0.7638 9.0000±3.5012 6.9833±3.1961 8.0843±3.3467 6.9321±3.2984
-CMAES 4.5833±2.2004 5.0000±3.6056 6.3333±2.0817 8.5833±3.1103 6.3167±2.1794 8.8313±3.7776 9.6605±4.5537
-Random Forest 6.6667±4.4121 8.3333±4.6188 9.5000±4.4441 11.7000±3.6237 12.5000±1.9343 8.8313±4.0817 8.7778±4.5600
-Gradient Boosting 10.5000±3.2711 15.6667±0.2887 14.5000±1.5000 8.4667±4.2729 12.4000±3.1139 9.1024±4.9840 9.0123±5.0812
-SVM 3.5000±3.3317 8.0000±1.7321 8.6667±5.1316 10.1167±4.8382 13.3667±2.6682 7.0602±3.9296 10.5185±5.4906
-Linear 7.8333±2.6394 11.6667±3.5119 11.3333±2.0817 5.7500±3.2450 7.5167±2.9870 7.7892±5.2179 9.0185±5.1705
-MA 12.8333±2.4014 4.6667±0.5774 2.3333±1.5275 10.0667±3.5349 11.3667±2.9271 8.4277±4.2949 8.7531±3.7200
-DivBO 9.4167±3.3229 10.6667±1.1547 12.6667±1.5275 10.0833±2.9916 7.3833±3.2262 8.1506±3.6599 8.4938±3.4228
-EO 10.0000±4.6476 4.6667±3.2146 4.0000±2.0000 9.3667±2.8555 8.5000±2.6425 8.1988±4.2677 7.6852±2.8410
-NE-Stack (Ours) 4.6667±3.0768 3.3333±1.1547 7.3333±0.5774 3.4500±2.3095 2.7167±2.2232 8.9639±5.4352 7.6173±5.5437
-NE-MA (Ours) 7.0000±1.7889 2.0000±1.0000 3.0000±2.0000 5.3500±3.7374 3.9667±2.2967 7.3012±3.9900 6.7037±3.7855
-Table 13: Average NLL for Subset of Base Models with DivBO
-Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
-Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
-Single DivBO 1.0000±0.0000 1.0000±0.0000 0.8707±0.3094 1.7584±2.0556 1.1846±0.2507 1.1033±0.9951 1.0039±0.0424
-Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9724±0.0478 1.1962±1.0189 0.9717±0.1919 1.0107±0.3431 1.0302±0.1250
-Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9762±0.0413 1.1631±0.9823 0.9431±0.2035 1.0023±0.3411 1.0247±0.1473
-Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9646±0.0614 1.2427±1.1130 0.9544±0.2050 1.0014±0.3423 1.0400±0.1949
-Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9762±0.0413 1.1659±0.9789 0.9435±0.2029 1.0024±0.3410 1.0271±0.1531
-MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9678±0.0558 1.0559±0.7452 0.9501±0.1617 1.0068±0.4141 1.0237±0.1502
-NE-Stack (Ours) DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.5040±1.9442 0.8329±0.2659 0.9729±0.3952 6.9453±3.4749
-NE-MA (Ours) DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 1.1237±0.9964 0.9200±0.1966 1.0016±0.3407 1.0070±0.0977
-Table 14: Average NLL for Subset of Base Models randomly
-Seletor FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
-Single - 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
-Single Random 1.0067±0.0164 1.0000±0.0000 0.9240±0.3504 1.2915±0.9952 1.1261±0.3134 1.0225±0.3353 1.1378±0.4641
-Top25 Random 0.8397±0.1000 0.5848±0.1980 0.6526±0.3019 3.6553±2.7053 3.0436±2.1378 1.2599±1.5015 1.0611±0.2799
-Quick Random 0.7305±0.0764 0.5958±0.1917 0.6656±0.2968 1.7769±2.1443 1.1646±0.3728 1.0797±1.0007 1.0151±0.1546
-Greedy Random 0.7024±0.0720 0.5783±0.1857 0.6617±0.2839 1.6723±2.1446 0.9961±0.1290 1.0725±0.9978 1.0023±0.0961
-MA Random 0.9069±0.1812 0.8677±0.2292 0.6698±0.2898 4.8593±3.1360 3.4575±2.6490 1.4759±1.9396 1.4286±1.7242
-NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6187±0.2950 0.8292±0.5466 0.8160±0.3852 0.9540±0.5077 4.2183±3.4808
-NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.6650±0.2750 1.6877±2.1535 1.0903±0.2578 1.0674±0.9998 1.0277±0.1994
-Table 15: Average NLL for Subset of Base Models (more baselines)
-Selector FTC NB (100) NB (1000) QT-Micro QT-Mini TR-Class TR-Reg
-Single DivBO 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000
-Random DivBO 0.9305±0.3286 0.6538±0.2123 0.9767±0.0404 0.8220±0.3507 0.8814±0.2062 0.9748±0.1415 1.0287±0.1365
-Top25 DivBO 0.7617±0.1136 0.5564±0.1961 0.9805±0.0339 0.7916±0.3245 0.8569±0.2178 0.9672±0.1381 1.0232±0.1550
-Quick DivBO 0.7235±0.0782 0.6137±0.1945 0.9687±0.0542 0.8048±0.3271 0.8672±0.2195 0.9652±0.1353 1.0378±0.1959
-Greedy DivBO 0.7024±0.0720 0.6839±0.3003 0.9805±0.0339 0.7910±0.3209 0.8572±0.2174 0.9673±0.1384 1.0256±0.1601
-CMAES DivBO 0.8915±0.1759 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 0.9671±0.1179 1.0250±0.1595
-Random Forest DivBO 0.7932±0.1194 0.9338±0.3436 1.1678±0.2906 1.6807±1.1160 1.5507±0.8449 1.3195±0.6428 1.0534±0.1277
-Gradient Boosting DivBO 0.7908±0.1848 1.4011±0.5359 1.0000±0.0000 2.4553±1.1191 2.7220±1.3007 1.2106±0.5937 1.1246±0.4052
-Linear DivBO 0.7433±0.0870 0.6471±0.2272 1.0248±0.0430 0.9690±0.4209 1.1132±0.4525 1.0814±1.0056 1.0316±0.1403
-SVM DivBO 0.8312±0.0943 0.7406±0.2612 1.0786±0.1362 5.2463±3.4452 4.8194±3.5169 1.5156±1.5597 2.7777±2.9458
-MA DivBO 0.7245±0.0788 0.5712±0.2185 0.9719±0.0486 0.7603±0.2876 0.8802±0.1741 0.9621±0.1306 1.0222±0.1570
-Single Random 1.0067±0.0164 1.0000±0.0000 1.0540±0.0936 0.8958±0.3297 0.9786±0.2444 0.9902±0.1529 1.1314±0.4475
-Top25 Random 0.8397±0.1000 0.5848±0.1980 0.7283±0.1534 2.7131±1.5510 2.5961±1.9153 1.2873±1.5773 1.0588±0.2762
-Quick Random 0.7305±0.0764 0.5958±0.1917 0.7467±0.1420 0.9847±0.1836 0.9737±0.1340 1.0043±0.2497 1.0123±0.1529
-Greedy Random 0.7024±0.0720 0.5783±0.1857 0.7478±0.1352 0.9171±0.0932 0.8960±0.0795 0.9762±0.1082 1.0002±0.1038
-CMAES Random 1.2164±0.3851 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.0000±0.0000 1.1544±0.9890 0.9985±0.0954
-RF Random 0.7505±0.0915 0.8325±0.2255 0.9850±0.0175 2.9838±1.7970 2.6001±1.8694 1.2006±0.4367 1.0058±0.0898
-GBT Random 0.7235±0.1605 1.8377±1.4510 1.2177±0.3771 1.9521±1.2076 2.8683±1.9763 1.2329±0.6934 1.0470±0.1913
-Linear Random 0.7541±0.0897 0.7400±0.2827 0.9103±0.1384 1.1776±0.3187 1.5498±0.8446 1.0417±1.0146 1.0408±0.1296
-SVM Random 0.8010±0.0903 0.7767±0.3006 0.9821±0.3676 5.1993±3.4269 5.1080±3.2902 1.4479±1.4990 2.8328±3.0214
-MA Random 0.9069±0.1812 0.8677±0.2292 0.7543±0.1312 4.1694±2.5844 3.0315±2.4578 1.5059±1.9654 1.4185±1.6833
-NE-Stack Random 0.7709±0.2204 0.7551±0.2493 0.6882±0.1602 0.6925±0.4510 0.6854±0.2801 0.9387±0.4628 4.2065±3.4678
-NE-MA Random 0.6972±0.0712 0.7911±0.2147 0.7535±0.1164 0.9291±0.0999 0.9191±0.0624 0.9922±0.2525 1.0259±0.2002
-NE-Stack DivBO 0.7715±0.2141 0.6204±0.2234 1.0000±0.0000 1.0442±0.8691 0.7482±0.2254 0.9671±0.3746 6.9574±3.4368
-NE-MA DivBO 0.7036±0.0698 0.5704±0.2345 1.0000±0.0000 0.7632±0.3319 0.8349±0.2075 0.9677±0.1455 1.0057±0.1121
-Figure 8: Average Normalized NLL in Scikit-learn pipelines (Validation).
+Table 11: Metadatasets Information
+Meta-Dataset Modality Task Information No. Datasets Avg. Samples for Validation Avg. Samples for Test Avg. Models per Dataset Avg. Classes per Dataset
+Nasbench (100) Vision NAS, Classification (Dong and Yang,, 2020) 3 11000 6000 100 76.6
+Nasbench (1K) Vision NAS, Classification (Dong and Yang,, 2020) 3 11000 6000 1K 76.6
+QuickTune (Micro) Vision Finetuning, Classification (Arango et al., 2024a, ) 30 160 160 255 20.
+QuickTune (Mini) Vision Finetuning, Classification (Arango et al., 2024a, ) 30 1088 1088 203 136.
+FTC Language Finetuning, Classification, Section (Arango et al., 2024b, ) 6 39751 29957 105 4.6
+TabRepo Clas. Tabular Classification  (Salinas and Erickson,, 2023) 83 1134 126 1530 3.4
+TabRepo Reg. Tabular Regression  (Salinas and Erickson,, 2023) 17 3054 3397 1530 -
+Sk-Learn Pipelines. Tabular Classification, Section O 69 1514 1514 500 5.08